### v1

# Multitask methods for predicting molecular properties from heterogeneous data

Abstract
Data generation remains a bottleneck in training surrogate models to predict molecular properties. We demonstrate that multitask Gaussian process regression overcomes this limitation by leveraging both expensive and cheap data sources. In particular, we consider training sets constructed from coupled-cluster (CC) and density function theory (DFT) data. We report that multitask surrogates can predict at CC level accuracy with a reduction to data generation cost by over an order of magnitude. Of note, our approach allows the training set to include DFT data generated by a heterogeneous mix of exchange-correlation functionals without imposing any artificial hierarchy on functional accuracy. More generally, the multitask framework can accommodate a wider range of training set structures—including full disparity between the different levels of fidelity—than existing kernel approaches based on $\Delta$-learning, though we show that the accuracy of the two approaches can be similar. Consequently, multitask regression can be a tool for reducing data generation costs even further by opportunistically exploiting existing data sources.
preprint: AIP/123-QED
I Introduction
Predicting molecular properties based on first-principle approaches drives innovation across the physical sciences. However, exact prediction is infeasible due to the exponential scaling of the Schrödinger equation—the underlying quantum-mechanical description of materials systems. Decades of research have yielded numerous first-principle models to describe electronic structures approximately, each representing a compromise between computational cost and prediction accuracy. Coupled cluster (CC) methods are typical examples of high fidelity models which offer accuracy in exchange for considerable computational cost. Specifically, CCSD(T)—the “gold standard” of quantum-chemical prediction—suffers from a steep $O(N^{7})$ scaling where $N$ is the number of electrons in the target system. Harding et al. (2008) In contrast, density-functional theory (DFT) methods are popular baseline models since they provide adequate accuracy for most systems while enjoying substantially lower cost (scaling only as $O(N^{3})$). Ying, Yu, and Ying (2019); Perdew and Schmidt (2001); Goerigkab and Grimme (2011); Teale et al. (2022)
Traditional workflows force users to settle on a single model which is cheap enough to result in manageable cost but still sufficiently high fidelity to give meaningful predictions. However, the advent of modern data-driven modeling has introduced new tools which make it possible to decouple expensive data generation from prediction. Statistical surrogates are trained with predictions from first-principle models and, once trained, can be substituted into workflows, making the prediction step considerably cheaper. In the context of building potential energy surfaces and interatomic potentials, these data-driven approaches have become the de facto standard. Smith et al. (2019); Dral et al. (2020); Goodlett, Turney, and Schaefer (2023); Zaverkin et al. (2023) Simultaneously, large-scale data generation is facilitated by sophisticated workflow engines for high-throughput calculations Curtarolo et al. (2012); Jain et al. (2011); Huber et al. (2020) and robust DFT methods inspired by mathematical research. Teale et al. (2022); Cancès et al. (2022); Herbst and Levitt (2020, 2022); Cancès et al. (2023) The increasing amount of openly available data Ruddigkeit et al. (2012); Ramakrishnan et al. (2014); Chanussot* et al. (2021); Smith, Isayev, and Roitberg (2017) could in principle be readily employed, but quality is generally not homogeneous as each dataset’s author has made different modeling choices. Consequently, most surrogate models are single fidelity—all training data is generated using the same quantum-chemical method and the same numerical approach. Thus, heterogeneity has been an obstacle to opportunistically exploiting the already available data. With this work, we contribute to the promising development of multifidelity surrogates which do not require a single choice between accurate (but expensive) and fast (but crude) methods, but instead combine the advantages of both. Specifically, we consider multitask methods which construct a relationship between multiple regression problems without requiring precise information about the respective accuracy of each source. Bonilla, Chai, and Williams (2008); Leen, Peltonen, and Kaski (2012)
Gaussian process (GP) regression is a popular tool for constructing surrogates informed by heterogeneous data. Several works Pilania, Gubernatis, and Lookman (2017); Batra et al. (2019) have explored the use of multifidelity GP models Kennedy and O’Hagan (2000) trained using two fidelities, namely a high and a low fidelity DFT approach. Alternatively, $\Delta$-learning may be employed to model the difference between two levels of theory, then by adding that difference to low level predictions of the target to produce high level predictions. Ramakrishnan et al. (2015); Dral, Zubatiuk, and Xue (2023); Batra et al. (2019) Though GP regression and the related technique of kernel ridge regression are common choices for fitting such models, in principle, any method of fitting a surrogate may be employed. Batra et al. (2019); Dral, Zubatiuk, and Xue (2023). $\Delta$-learning can be extended to leverage more than two training data sets if multiple difference models are learned and summed. Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) However, between each pair of training sets, data points must be in some way aligned to allow for learning a difference model. For example, in a model of the difference between DFT and CCSD(T), DFT predictions may only contribute to the training set if a corresponding CCSD(T) prediction is also available. Further, when either $\Delta$-learning or multifidelity GP regression learn from more than two heterogeneous training sets, these sets must be assigned some order—for example, following a hierarchy in accuracy. Recent work Vinod, Kleinekathöfer, and Zaspel (2023) has proposed an optimized linear combination of submodels as an alternative, but all coefficients in this linear combination must be re-optimized whenever any submodel is altered, and opportunistic data set construction is not fully explored.
Progress towards multifidelity approaches has also been made in the context of neural network surrogates. For example, transfer learning methods train an initial neural network on a larger source dataset, then re-optimise part of the parameters using a target dataset. Towell and Shavlik (1994); Fu (1989); Smith et al. (2019) For the prediction of atomistic properties, the source dataset might contain millions of DFT calculations while the target might be a smaller set of CCSD(T) predictions. In prior studies, it has been observed that transfer learning can produce higher accuracy than training a network on only one of these datasets. Dral et al. (2020); Dral, Zubatiuk, and Xue (2023) While the framework we discuss in this work is not restricted to GP regression strategies, we will focus exclusively on GP methods. In comparison to neural network approaches, where a rigorous mathematical explanation for the low generalization error of well-trained models is still lacking Bartlett, Montanari, and Rakhlin (2021); Lotfi et al. (2022), GP approaches are more amenable to mathematical analysis. Rasmussen and Williams (2006) This comparative transparency simplifies interpretation of the surrogate’s predictive process in the diverse training data configurations we analyze here. Additionally, GP methods have the advantage of yielding accurate predictions even for relatively small training set sizes. While the training cost of GP regression may become large as the size of the training set increases, many strategies for efficient scaling have been developed and can be routinely employed Quinonero-Candela and Rasmussen (2005); Liu et al. (2018); Wilson, Dann, and Nickisch (2015); Cole, Christianson, and Gramacy (2021).
This work explores a GP regression approach which can accommodate an arbitrary number of quantum chemical datasets without an assumption of hierarchy. In particular, we employ the multitask framework Leen, Peltonen, and Kaski (2012), which has been shown to easily incorporate heterogeneous data sources. The notion of heterogeneity in fact encompasses many possibilities. Fig. 1 compares the multitask method to other approaches by visualizing mappings between sets of molecules and the quantum chemistry methods which are used to generate viable training data sets. The highest level of theory used to construct these training sets will be called the primary level. In our case, the primary level will be CCSD(T). All other levels of theory will be categorized as secondary levels. Our secondary levels will include DFT methods using different density functional approximations (DFAs). Different levels of theory may contribute data points corresponding to the same set of molecules, different sets of molecules, or some union of shared and unique sets of molecules. Furthermore, the testing data set contains primary level predictions for target molecules. Thus, the multitask method has the option to use secondary predictions for the target molecules as part of the training set. The diagrams in Fig. 1 show several options for constructing a training set for inference. The impact of these choices on inference is under-explored in literature, and the multitask method is a useful tool for investigating the benefits of different training set configurations. Crucially, the multitask framework enables training with a “dataset of opportunity,” amalgamated from existing data sets, as an alternative to expending computational time to generate entirely new data.
Figure 1: Diagram of training data sets compatible with the methods considered in this work. Here, we define three sets of molecules—Core, Additional, and Target systems—and two quantum chemistry methods—primary and secondary. Our goal is to train a model that will predict our quantity of interest for the Target molecules with the accuracy of the primary method. By definition, the Core molecules are all systems for which we can compute training data with the primary method. Additional molecules are systems for which we have no primary data but which are not a part of the Target set. Theoretically, we may train using secondary predictions for molecules of all sets though not all inference methods support use of all data sets. The diagram shows the combinations of data sets and quantum chemistry methods supported by each inference method.
I.1 Advantages of the Multitask Method
Our numerical experiments will demonstrate the following advantages of the multitask modeling approach. Key figures illustrating each claim are referenced in parentheses.
  • Multitask models can achieve target accuracy with less computational cost than Gaussian process regression (Fig. 3).
  • Accuracy of multitask inference improves as more DFT data is added to the training set. Furthermore, increasing the amount of training data from different secondary methods for the same molecule improves accuracy (Fig. 6).
  • Unlike the $\Delta$ and multifidelity approaches, multitask methods can model more than two observation data sets without imposing an arbitrary order on the sets (Figs. 5 and 6)
  • Multitask models have flexible training set requirements. The sets are constructed from multiple levels of theory, and the same molecules do not need to be included for each level (Fig. 3).
Flexibility is a key advantage of the multitask method compared to a single task or $\Delta$-learning approach. The method is compatible with many training set structures. For instance, a multitask model can be trained with CCSD(T) data for one molecule set and DFT data for a different molecule set, with CCSD(T) and DFT data for the same molecule set, or with some in intermediary case. Models also have the option to train with DFT data for the molecules that we aim to make predictions for at the CCSD(T) level of accuracy. The flexible multitask framework makes it possible to explore the impact of different compositions of a training data set on inference as seen in Fig. 3. Consequently, we can determine the balance of expensive high level and cost effective low level data which best suits the problem setting and available data.
In Section II, we will review prediction with GP regression as well as the $\Delta$ and multifidelity methods before describing the multitask framework. Section III provides detail on our approach to testing different training set configurations, and Section IV describes our numerical tests. We demonstrate the performance of the multitask method on two examples: prediction of three-body energy of water trimer configurations and prediction of ionization potential of small organic molecules.
II Statistical Models
We review the formulation of Gaussian process regression models. Subsections II.2 through II.4 discuss variants of GP regression which leverage heterogeneous sets. See references Rasmussen and Williams (2006); Kennedy and O’Hagan (2000); Leen, Peltonen, and Kaski (2012); Bonilla, Chai, and Williams (2008); Abrahamsen (1997) for more detail.
II.1 Gaussian Process Regression
Our goal is to predict values of some quantity of interest (QoI) $f$, given an identifying input feature $\bm{X}$. We train our predictive model using $\{\bm{X}_{i},\bm{Y}_{i}\}_{i=1}^{n}$ where each $\bm{Y}_{i}$ is a noisy observation of $f(\bm{X}_{i})$. In our setting, we interpret imperfect first-principle predictions of the QoI as noisy observations of the true quantity. Each $\bm{X}_{i}$ is some representation of the electronic structure of a molecular system. For our numerical results, we build $\bm{X}_{i}$ using the well established Smooth Overlap of Atomic Potentials (SOAP). Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) $\bm{Y}_{i}$ is the CCSD(T) or DFT prediction of the quantity of interest for the $i^{th}$ molecular configuration.
We will assume that each observation is independently perturbed by additive noise which follows identical Gaussian distributions. The corresponding model is
$\bm{Y}_{i}$ $=$ $\ f(\bm{X}_{i})\ +\ \varepsilon_{i},$ (1)
$\varepsilon_{i}$ $\sim$ $\ N\left(0,\ \sigma^{2}\right).$
Before inference, we choose a mean function, $\mu(\bm{X})$ to describe the expected value of $f(\bm{X})$ and a kernel function, $k(\cdot,\cdot)$, to encapsulate our assumptions about the relationship between $f(X_{i})$ and $f(X_{j})$. Typically, $k(\cdot,\cdot)$ will have hyperparameters that are selected through optimization, then fixed during inference.
We will employ a Gaussian process prior model for our regression function:
$f(\bm{X})\ \sim\ \text{GP}\Big{(}\mu(\bm{X}),k(\bm{X},\bm{X}^{\prime})\Big{)}.$ (2)
As a consequence of this prior assumption, any finite collection, $\{f(\bm{X}_{1}),\dots,f(\bm{X}_{n}\})$, will follow a multivariate normal distribution with mean $\bm{\mu}=\{\mu(\bm{X}_{1}),\dots,\mu(\bm{X}_{n})\}$ and covariance matrix $K$ with entries at row $i$ and column $j$ given by $k(\bm{X}_{i},\bm{X}_{j})$. We are interested in inferring the values the function takes on at target inputs, and we collect these values in the vector, $\bm{f}_{*}$. Using this notation, we write the joint distribution on the observations and targets for inference:
$\begin{bmatrix}\bm{Y}\\ \bm{f}_{*}\end{bmatrix}\ \sim\ N\left(\begin{bmatrix}\bm{\mu}\\ \bm{\mu}_{*}\end{bmatrix},\ \begin{bmatrix}K_{pp}+\sigma^{2}I&K_{p*}\\[3.0pt] K_{p*}^{T}&K_{**}\end{bmatrix}\right).$ (9)
where $I$ is the identity matrix. The specific subscripts of each covariance matrix represent the features that the matrix compares: $p$ corresponds to our primary training data and *** indicates the target set. We note that in this single-fidelity setting all training data is contained in the primary set, but this will no longer be the case when we consider surrogates informed by heterogeneous data.
The marginal distribution of $\bm{Y}$ is the model evidence, and conditioning $\bm{Y}$ on $\bm{f}_{*}$ yields the likelihood. To obtain the posterior, we employ the well known formula for conditioning part of a multivariate Gaussian random variable on the rest, Rasmussen and Williams (2006) finding that
$\bm{f}_{*}\ \big{|}\bm{Y}\ \sim\ N\Big{(}\bm{\mu}_{*}\ +\ K_{p*}^{T}(K_{pp}+\sigma^{2}I)^{-1}\big{(}\bm{Y}-\bm{\mu}\big{)},$ (10)
$\quad K_{**}\ -\ K_{p*}^{T}(K_{pp}+\sigma^{2}I)^{-1}K_{p*}\Big{)}.$
Eq. (10) gives the probability distribution of our target values, conditioned on our observations. The mean of this distribution is a correction to our prior mean based on the deviations of the observations from that mean. The predicted covariance indicates how much our observations have reduced our uncertainty, provided our assumptions are reasonably aligned to reality.
II.2 $\Delta$-Learning
Suppose that for each training input, we have data on our QoI from multiple sources. We thus seek an inference model that leverages $\left\{\bm{Y}_{ij}\right\}_{i=1,j=1}^{n,m}$, where we have data for molecular configurations $i=1,\dots,n$ and methods $j=1,\dots,m$. A simple “pooling” approach may put all observations into one vector and proceed with inference via Eq. (10). Then, $\bm{Y}_{ij_{1}}$ and $\bm{Y}_{ij_{2}}$ are interpreted as independent observations of the QoI, $f(\bm{X}_{i})$, for molecular configuration $\bm{X}_{i}$. The distinct behavior of each quantum chemical method is treated as a perturbation of some mean value by a realization of the noise distribution. We seek instead a model where the distribution of each $\bm{Y}_{ij}$ is sensitive to $j$. Ideally, we want to map differences in electronic structures methods to differences in their predictions.
Method-specific features could be a solution, but construction of such features is challenging. One approach involves appending a vector which indicates the method category to each feature. This strategy puts a large burden on covariance hyperparameter optimization since one set of hyperparameters must encode differences between molecular configurations and between quantum chemical methods. Alternatively, to characterize the qualitative difference between methods within the features, we would have to solve the very problem for which we are creating the features in the first place—for instance, we would need to reliably predict CCSD(T) performance given CCSD(T) and DFT data from training molecules as well as DFT data for target molecules. Instead, we use this subsection to discuss $\Delta$-learning, which directly models the differences between the predictions of quantum chemistry methods.
$\Delta$-learning performs Gaussian process regression to predict differences, $\Delta_{i}$, between observations obtained with two different methods for a given $\bm{X}_{i}$. If we retain our assumption that noise follows an independent, identically distributed Gaussian distribution, we have the model
$\bm{Y}_{ij_{1}}-\bm{Y}_{ij_{2}}$$\equiv$ $\ \Delta_{i}\ =\ f\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{i},$ (11)
$f$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}_{\Delta},\ k_{\Delta}(\bm{X},\bm{X}^{\prime})\big{)}.$
where $j_{1}$ and $j_{2}$ indicate two prediction methods. An application of Eq. (10) yields a distribution on the regression function for $\Delta$. Suppose our ultimate interest is to predict $\bm{Y}$ at $\bm{X}_{*}$. Within the $\Delta$ framework, we can supply an observation from one method, say $\bm{Y}_{*j_{2}}$, in order to make a prediction for the value from the other method:
$\bm{Y}_{*j_{1}}\ \approx\ \Delta_{i}\ +\ \bm{Y}_{*j_{2}}.$
Thus, $\Delta$-learning is suited to settings where calculations by method $j_{1}$ are both more accurate and more time consuming than calculations that use $j_{2}$.
Consider a case where we want to build a model from an existing data set which includes predictions from CCSD(T) and several DFT methods. If the predictions all correspond to the same set of molecules, $\Delta$-learning models have the ability to leverage more of the data set than conventional GP regression models. Unfortunately, $\Delta$-learning is not-applicable in cases where predictions by different methods correspond to different molecular configurations. To use a DFT prediction for a molecule $\bm{X}_{i}$ to train a $\Delta$ model between CCSD(T) and DFT, we also require a CCSD(T) prediction for the same molecule. In the following subsections, we consider alternative approaches which model disparities in data sources and have the ability to use all available observations.
II.3 Multifidelity Method
As a first step toward considering statistical models which train on heterogeneous data, we describe the multifidelity data fusion model. Kennedy and O’Hagan (2000) The multifidelity approach defines two regression functions: $f_{p}$,representing the predictions from some high fidelity model, and $f_{s}$, representing a low fidelity model. Kennedy and O’Hagan (2000) We assume the relationships
$\bm{Y}_{ip}$ $=$ $\ f_{p}\big{(}\bm{X}_{i}\big{)}\ +\varepsilon_{ip}$ (12)
$=$ $\ \rho f_{s}\big{(}\bm{X}_{i}\big{)}\ +\ \delta_{ps}\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{ip},$
$\bm{Y}_{is}$ $=$ $\ f_{s}\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{is}.$ (13)
The term $\delta_{ps}$ captures the difference between the high fidelity model and the low fidelity model, scaled by hyperparameter, $\rho$. This disparity is endowed with a Gaussian process prior, as is the low fidelity regression function, $f_{s}$. The correlation, $\rho$, will be optimized before inference. As before, we assume that the additive noise terms, $\varepsilon_{ip}$ and $\varepsilon_{is}$, are independent and drawn from identical, centered Gaussian distributions. Now, suppose we want to predict our targets, $\bm{f}_{*}$, at the high fidelity level. We can determine that the joint distribution of our high and low fidelity observations and our targets has covariance
$\Sigma\ =\ \begin{bmatrix}\rho^{2}K^{s}_{pp}+K^{\delta}_{pp}+\sigma^{2}I&\rho K^{s}_{ps}&\rho^{2}K^{s}_{p*}+K^{\delta}_{p*}\\[5.0pt] \rho K^{s}_{sp}&K^{s}_{ss}+\sigma^{2}I&\rho K^{s}_{s*}\\[5.0pt] \rho^{2}K^{s}_{*p}+K^{\delta}_{*p}&\rho K^{s}_{*s}&\rho^{2}K^{s}_{**}+K^{\delta}_{**}\\[5.0pt] \end{bmatrix}.$ (17)
where $\sigma^{2}$ is the variance of the noise distribution. The superscript of each covariance matrix, K, indicates which Gaussian process prior corresponds to the kernel function used in its construction, and the subscripts indicate the pairs of features that are compared. Bayesian inference yields an analytical posterior distribution, analogous to Eq. (10).
II.4 Multitask Method
“Multitasking” refers to a category of methods which consider several Gaussian process regression tasks and assume some relationship between these tasks. As with $\Delta$-learning or multifidelity fusion, multitasking can allow us to incorporate a larger dataset into our inference problem without just pooling the data into one observation vector. For instance, we can define a regression problem trained on CCSD(T) data as well as one trained on each DFA we consider. By modeling correlation between regression functions, we can use the DFA regression tasks to support prediction in the CCSD(T) task.
We can apply multitask regression to problems with $m$ tasks where, for each task $j\in\{1,\ldots,m\}$, we have observational data $\bm{Y}_{j}\in\mathbb{R}^{n}$. In practice, the method can be generalized, so that observation sets of different sizes are used for different tasks. Symmetric variants of multitask modeling treat all tasks equally. For instance, we may assume that all the data are modeled by a single Gaussian process whose covariance kernel is the Kronecker product of one kernel which relates the tasks and another which models relationships between data from the same task.Bonilla, Chai, and Williams (2008)
In many cases, however, the goal is to perform regression for a primary task, and all other tasks may be considered secondary tasks which provide data that may be useful in learning the primary task. For our own setting, CCSD(T) is a natural choice of data for the primary regression task because we would like to make predictions which match this method in accuracy. Each set of DFT predictions made using a different DFA could inform a secondary regression task which supports the primary task. Thus, we consider an asymmetrical model Leen, Peltonen, and Kaski (2012) in which secondary tasks are related through the primary task.
Suppose $\bm{Y}_{p}\in\mathbb{R}^{n_{p}}$ is the observation data for the primary task and $\bm{Y}_{s_{j}}\in\mathbb{R}^{n_{j}}$ is the data for the $s_{j}^{th}$ secondary task. As before, we assume that each task has its own regression function
$\bm{Y}_{ip}$ $=$ $\ f_{p}(\bm{X}_{i})\ +\ \varepsilon_{ip},$ (18)
$\bm{Y}_{is_{j}}$ $=$ $\ f_{s_{j}}(\bm{X}_{i})\ +\ \varepsilon_{is_{j}}\qquad\forall j=1,\dots,m.$ (19)
where each noise term is drawn independently from a centered Gaussian distribution with variance $\sigma^{2}$. We now model a shared structure in the regression functions based on the primary function:
$f_{s_{j}}(\bm{X}_{i})$ $=\ \rho_{s_{j}}f_{p}(\bm{X}_{i})\ +\ \delta_{s_{j}}(\bm{X}_{i})\qquad\forall j=1,\dots,m.$ (20)
The specific component, $\delta_{s_{j}}$, of the secondary function serves to “explain away” behavior that is not captured by the shared component, $f_{p}$een, Peltonen, and Kaski (2012) Both the correlation parameter, $\rho_{s_{j}}$ and the specific component aim to mitigate negative transfer—learning behavior from secondary tasks that is not representative of the primary task. We make the prior assumption that $f_{p}$ and $\delta_{s_{j}}$ for $j=1,\dots,m$ follow Gaussian process distributions, each with a mean and kernel function which we will specify.
As was the case in subsections II.1 through II.3, the first step toward inference with the multitask method is to determine the joint distribution on observations and targets. Note that the specific component of each task is modeled as independent of the others. Then, the covariance between observations due to the specific component is a block matrix with zeros corresponding to the covariance of specific functions for different tasks. Call this block matrix $K^{\delta}$. Let $\bm{R}$ be a diagonal matrix with dimensions equal to the number of observations, and let each diagonal element be the correlation hyperparameter of the corresponding task. The covariance of the joint distribution is
$\Sigma\ =\ \begin{bmatrix}K^{p}_{pp}+\sigma^{2}I&K^{p}_{ps}\bm{R}&K^{p}_{p*}\\[5.0pt] \bm{R}K^{p}_{sp}&\bm{R}K^{p}_{ss}\bm{R}+K^{\delta}+\sigma^{2}I&\bm{R}K^{p}_{s*}\\[5.0pt] K^{p}_{*p}&K^{p}_{*s}\bm{R}&K^{p}_{**}\\[5.0pt] \end{bmatrix}.$ (24)
The superscript $p$ indicates that a matrix was created with the primary kernel, and the matrix subscripts indicate the features that are compared. Inference proceeds as described in Eq. (10).
II.5 Beyond Two Tasks
Asymmetric multitasking assumes a relationship between regression functions—given by Eq. (20)—which is structurally reminiscent of the multifidelity fusion model—Eq. (12). Both assume that the regression function of one dataset is related to the regression function of another by a scaling parameter, $\rho$, and an additive disparity function, $\delta$.
For only one secondary model, $s_{1}$, the relationships between the regression functions posited by both methods are mathematically equivalent. However, asymmetric multitasking diverges more significantly from the multifidelity method when more than two models are considered. In the multifidelity approach, model hierarchy is reflected in a nested relationship between the functions,
$f_{p}(\bm{X})$ $=$ $\ \rho_{s_{1}}f_{s_{1}}(\bm{X})\ +\ \delta_{s_{1}}(\bm{X}),$
$f_{s_{1}}(\bm{X})$ $=$ $\ \rho_{s_{2}}f_{s_{2}}(\bm{X})\ +\ \delta_{s_{2}}(\bm{X}),$
$\vdots$
$f_{s_{j-1}}(\bm{X})$ $=$ $\ \rho_{s_{j}}f_{s_{j}}(\bm{X})\ +\ \delta_{s_{j}}(\bm{X}),$
whereas in the asymmetric approach, all functions share the same relationship to the primary function
$f_{s_{1}}(\bm{X})$ $=$ $\ \rho_{s_{1}}f_{p}(\bm{X})\ +\ \delta_{s_{1}}(\bm{X}),$
$\vdots$
$f_{s_{j}}(\bm{X})$ $=$ $\ \rho_{s_{j}}f_{p}(\bm{X})\ +\ \delta_{s_{j}}(\bm{X}).$
Thus, compared to multifidelity fusion, the asymmetric multitasking approach more naturally accommodates a reasonably large number of regression tasks and relates each closely to the primary task of interest. Multitasking also holds a similar advantage over $\Delta$-learning. Before the notion of a difference can be imposed it is necessary to choose some order between the source data sets. Such an order may be arbitrary, for example, if our data sets are generated by multiple DFAs with similar structure. In Section III, we will describe different compositions of a training data set which the multitask framework allows us to test.
III Heterogeneous Training Data
In our numerical examples, we train inference models using data for multiple molecular configurations computed with multiple quantum chemistry methods. The multitask method leaves many questions about the particulars of the training data set up to the researcher. For instance, available data must be apportioned to different tasks. One of our goals is to demonstrate that by training with data from multiple quantum chemistry methods, we can construct an inference model with prediction accuracy comparable to CCSD(T) with less computational expense than if we trained a single task model on CCSD(T) alone. Thus, our CCSD(T) data is the primary target of our regression problem and a natural choice to inform the primary task. We let each secondary task correspond to a different DFA. Examples with one up to four secondary tasks will be considered.
Another decision to be made is whether the training data sets for different tasks should include the same molecular configurations. In other words, we must determine whether or not to train with predictions for the same quantity by different methods. If our target for inference is the CCSD(T) prediction of energy for a system of water molecules, we may also choose to include DFT predictions of this energy in the training set for a secondary task. The relative benefits of constructing training sets with different degrees of intersection between tasks is under-explored in literature. Our work seeks to understand the impact of training set design on model performance.
We consider a range of possible training set compositions here, and Fig. 2 provides two example diagrams of our system for distinguishing between different constructions. We assign the molecular configurations to three sub data sets—Core (C), Additional (A), and Target (T). Dots in the figure indicate when configurations are available at a given level of theory. Black indicates availability to the training set, and green indicates membership of the testing set. The C set contains all configurations for which CCSD(T) training data is available, the A set contains configurations for which we have DFT data but no CCSD(T) data, and the T set contains configurations that we use to test our model’s ability to reproduce CCSD(T) level accuracy.
Figure 2: Example data set structures. C, A, and T are sets of molecular configurations. C contains all configurations that the primary task, CCSD(T), is trained on. T contains the Target configurations; our goal is to estimate their properties with the accuracy of CCSD(T). The testing set is colored green. All secondary tasks are highlighted in gray. Configurations in set A can only be used for training secondary tasks. The vertical alignment of dots indicates that multiple tasks have access to the same molecular configurations.
In our numerical results, we always train models with the CCSD(T) data from the C set and consider various compositions of the secondary sets. All secondary tasks are highlighted in gray. We identify different training set compositions by the letters of the sets of molecule which are included in our secondary training data. For example, the top diagram in Fig. 2 shows an A training set because the only DFT predictions available to train the model come from the Additional set. By contrast, the lower diagram shows a CAT training set because dots are present in the slots for DFT predictions for Core, Additional, and Target molecular configurations.
Furthermore, the vertical alignment of dots indicate when some of the same molecular configurations are present in the training sets for multiple tasks. In the CAT example in Fig. 2, the dots indicate that the secondary tasks share $20\%$ of the C and A sets while the rest of the molecules in these sets are assigned to only one task each. In Section IV, we will show results for the C, CT, A, AT, CA, and CAT training set configurations. Fig. 6 shows the impact of varying the fraction of molecules that the secondary tasks have in common.
Our testing framework is designed so that models trained with DFT Target data can readily be compared with equivalent models that do not train on this data. Since the inference model can make Target predictions one at a time, for each Target molecule we train a new model using the C set, A set, and the DFT prediction for only that Target molecule. While in practice we could take the much more efficient approach of training on the entire DFT Target set at once, our approach enables us to contrast each CAT model to a CA model with a training set which only differs by one molecule. Consequently, we can isolate the effect of the DFT level Target data on prediction performance.
In Section IV, we will examine the performance of the $\Delta$ and Multitask methods, informed by CCSD(T) and DFT predictions, and compare them to basic Gaussian process regression, informed only by CCSD(T).
IV Numerical Examples
IV.1 Numerical Setup
We consider two case studies. In the first, the quantity of interest is the three-body (3-b) interaction energy of randomly selected water trimer configurations. The 3-b interaction term is isolated by subtraction of all 1-b and 2-b energies from the full energy of the trimer. We selected water because it is a simple molecule with unique properties which make it crucial for life and a popular subject for research. Chen et al. (2017); Dasgupta et al. (2021); Gillan, Alfè, and Michaelides (2016) The second case study considers the prediction of the ionization potential of small organic molecules. Ionization potential is known to be challenging for DFT to capture accurately, so it is insightful to test the ability of inference models to predict this quantity.
IV.1.1 Data Generation
For our first test case—inference of 3-b interaction energy—5986 water trimer configurations and corresponding CCSD(T) calculations were randomly selected from 45,332 configurations available from https://github.com/jmbowma/q-AQUAYu et al. (2022) Density functional theory data was generated for 2992 of these configurations using Psi4. Smith et al. (2020) We employed the PBEPerdew, Burke, and Ernzerhof (1996) and SCANSun, Ruzsinszky, and Perdew (2015) functional approximations. The former was selected for its popularity in materials science research and the latter because its description of intermediate range dispersion has been found to contribute to accurate prediction of water energy differences Chen et al. (2017); Dasgupta et al. (2021) Following recent work on dispersion corrections, Price, Bryenton, and Johnson (2021) we paired the PBE approximation with a D3 correction using finite damping and the SCAN approximation with a D3 correction using zero damping.  Both DFT and CCSD(T) calculations employed a standard counterpoise correction to alleviate basis set superposition error. Huang, Braams, and Bowman (2006)
The second test case is informed by the ionization potential of 3,165 molecular configurations, combinations of 479 different molecules and 7 configuration options. Duan, Fisher, and Kulik (2023) Configurations are subselected Duan et al. (2020) from the ANI-1 dataset Smith, Isayev, and Roitberg (2017) and contain only elements from the set $\{H,C,N,O\}$. We compute Duan et al. (2021) ionization potential for these configurations using CCSD(T) as well as secondary tasks corresponding to data from four density functionals: PBE0_DH, Brémond and Adamo (2011) PBE0, Adamo and Barone (1999) PBE, Perdew, Burke, and Ernzerhof (1996) and BLYP. Lee, Yang, and Parr (1988)
Quantification of cost in our numerical results refers to the cost of data set generation. The cost of inference is negligible next to quantum chemistry calculations. We base our cost model for each method on the average runtime of Psi4 computations for ten representative systems. For more details about the cost model, see Appendix A.
IV.1.2 Features
We require features which can distinguish between molecular systems to serve as inputs, $\bm{X}_{i}$, to our Gaussian process models. Kernel functions use features to model the relationships between quantities of interest. As noted in Section II.1, in our experiments we consider Smooth Overlap of Atomic Positions (SOAP) features, which have been successfully used to fit interatomic potentials in a Gaussian process based approach.  Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) SOAP features are constructed based on the neighborhood of each atom in a system, so challenges arise when two systems contain different numbers of constituent atoms of elements. Multiple strategies have been proposed for constructing “global” features to capture entire systems. De et al. (2016) The simplest approach is to take an average over the SOAP features for each atom in the system. Since this approach is computationally efficient and works well in practice, we apply it in our test cases. We chose SOAP parameters following experimentation and the advice of literature.  Deringer et al. (2021); Musil et al. (2021) More details are available in Appendices B and C.
IV.1.3 Hyperparameter Optimization
Squared exponential kernels are used within all Gaussian process priors in our experiments. Abrahamsen (1997) These kernels introduce mean, variance ($v$), and length scale ($\ell$) hyperparameters in addition to the multitask correlation ($\rho$). We use the mean of our training data for each task as that task’s prior mean, calculated at training time. To estimate the remaining parameters, we construct an optimization data set containing $150$ randomly selected molecular configurations which we subsequently exclude from our training and testing data sets. We assume here that we have access to predictions of our quantities of interest from each quantum chemistry method for each of these $150$ molecular configurations, but when this assumption does not hold, a conventional GP regression model may be used to fill in missing data.
We estimate correlation parameters, $\rho$, using Pearson’s correlation coefficient between each secondary task optimization data set and the corresponding primary task data. Popular belief in literature holds that correlation parameters require a detailed optimization procedure, Dral, Zubatiuk, and Xue (2023) but we find that these easy to obtain estimates work reasonably well in practice. A more detailed optimization is left for future work. We estimate $\ell\in\mathbb{R}$ by maximizing its log-marginal likelihood. This approach is easily extended to multifidelity and multitask cases by treating $Y_{ij}-\rho_{j}Y_{i1}$ as the $i^{th}$ observation. Forrester and A. Sóbester (2007) For the organic molecules example, we also use a maximum likelihood estimate for $v$ while for the water trimer example, we find better results from using an estimate of $v$ which minimizes mean absolute prediction error.
IV.2 Three-Body Interaction Energy of Water Trimer
IV.2.1 Multitask Performance
Figure 3: Water Trimer Case. Comparison of the MAE of the multitask method for a range of training set constructions (scatter points) to the MAE of a GP regression model trained only with CCSD(T) data (black line). Predictions by the multitask method are informed by CCSD(T) and SCAN. The legends of the plots indicate the structure of the secondary training set according to the system presented by Fig. 2. All MAE reported are the average results obtained from six randomly chosen training sets.
We first compare the performance of multitask models with only one secondary task to the performance of a traditional GP—or single task—model. Fig. 3 shows the error incurred by several implementations of each model when predicting the three-body (3-b) energy of water trimer configurations. We used the SCAN functional to generate the secondary task data, and Pearson’s correlation coefficient between the primary and secondary tasks is $0.997$, indicating strong positive correlation. The three subplots of Fig. 3 cover six secondary training set configurations (C, CT, A, AT, CA, and CAT), labeled within the legends.
All models were tested on a Target set of 320 molecular configurations. We let the size of the Core set, $N_{c}$, take on values in $\{5,10,20,40,80,160,320\}$ for both inference methods and additionally consider $N_{c}\in\{640,1280,2560\}$ for the single task method. The subplots of Fig. 3 also show seven clear groupings of data points with steep slope. This effect is a consequence of the high cost of CCSD(T) relative to DFT. Each of the seven groupings corresponds to a different $N_{c}$. Given the size of the Core set, we let the number of configurations in the A set range from $0.5N_{c}$ to $6N_{c}$. Each point on the plot represents a model with a different training set size. The x axis indicates the cost of the training set while the y axis gives the average mean absolute error for six random assignments of molecular configurations to the C, A, and T sets. We provide details of our cost model in Appendix A.
Refer to caption Refer to caption
(a) (b)
Figure 4: Water Trimer Case. Effect of including Core and Additional data in secondary training sets. Each indigo point represents a comparison between two multitask models where one model uses a training set which is a subset of the other. In (a), this comparison is between A and CA training configurations, and in (b), we compare AT and CAT configurations. For each pair of models compared, both the primary training data and the secondary data from the A set are identical. The $y=x$ line is plotted in green for reference.
For every training set configuration except C, the multitask method outperforms the single task reference, shown by a black line. Including T data from DFT in the training set leads to better results than corresponding cases without T in every example. In particular, training with secondary T data can enable the multitask method to exhibit an accuracy comparable to a GP model that is an order of magnitude more expensive.
While the C results demonstrate no significant improvement on the single task model, both A and CA results show steady improvement as the size of the A set increases. This behavior suggests that for the multitask method to be beneficial, secondary training data must cover molecular space that is not included in the CCSD(T) training data. We mention that it is possible that the behavior of C models would improve if we consider task dependent features or different relationships between task regression functions. We do not investigate these alternatives further here; instead, we focus on the usual setting where secondary DFT data for additional systems is cheap to obtain.
CT and CAT models are the best-performing multitask models, demonstrating very similar MAE per cost to each other. Furthermore, unlike the AT model, the CAT model does not show improvement as the size of the A set increases. It seems that for our setting, the combination of secondary C and T training data is so valuable to the model that any benefit from A data is dwarfed. We can hypothesize that when the model has access to data for the Core molecules from the primary and secondary tasks, it implicitly learns information about the difference between the tasks. Given secondary data on the Target, it can leverage the captured relationship between tasks to make an accurate prediction for the Target on the primary level.
Fig. 4 provides more insight to the relative roles of the training data sets and to the inherent flexibility of the multitask framework. While Fig. 3 demonstrates that multitask models can be successful when trained on many different training data configurations from the CAT framework, it does not directly compare the performance of the A and CA models. Such a comparison will illuminate the resilience of the multitask model to the loss of points in the training set which are common to primary and secondary tasks. We emphasize that this resilience is an advantage of the multitask method in comparison with $\Delta$-learning—where no training is feasible without access to secondary C data. Each point on subplot (a) of Fig. 4 compares the average MAE of two multitask models: one with a CA secondary training set, the other with an identical A set but no secondary C data. As before, we are inferring the 3-b energy of water trimer configurations with only one secondary task. The closeness of the points to the reference $y=x$ line shows that there is minimal effect on accuracy when the secondary C data is excluded from the training set. In fact, the median improvement in accuracy from using the CA training model instead of an A training model model is $0.0007$ eV. By comparison, the median improvement in accuracy from combining DFT data from the A set with CCSD(T) training data—that is, switching from a single task approach to a multitask method trained on an A set—is $0.004$ eV.
Subplot (b) of Fig. 4 also compares models trained with and without C secondary data in the case where all of these models also train on secondary Target data. While the AT models perform well, there is a penalty to losing access to secondary C training data when secondary Target data is also available. The median improvement in accuracy when switching from an AT model to a CAT model is $0.002$ eV, a third of the $0.006$ eV median improvement incurred from switching from a single task approach to an AT multitask model with the same CCSD(T) training cost. As we remarked previously, when the model trains on secondary C data, it likely gains information about the difference between the primary and secondary tasks. This information is most useful to the model when it also has access to secondary predictions for the Target molecules.
IV.2.2 Comparison with $\Delta$-learning
CCSD(T) [eV] SCAN [eV] PBE [eV]
absolute mean $0.0205$ $0.0223$ $0.0225$
absolute median $0.0126$ $0.0144$ $0.0155$
variance $0.0009$ $0.0010$ $0.0011$
CCSD(T)--SCAN [eV] CCSD(T)--PBE [eV]
absolute mean $0.0021$ $0.0037$
absolute median $0.0013$ $0.0018$
variance $1.1\times 10^{-5}$ $3.2\times 10^{-5}$
Table 1: Water Trimer case. Summary statistics for absolute quantum chemistry predictions and for differences.
(a) Estimation of CCSD(T)-PBE
(b) Estimation of CCSD(T)-SCAN
Figure 5: Water Trimer Case. The MAE of different inference methods: GP regression (line), multitask regression trained with a CAT secondary set (orange), two level $\Delta$-learning (indigo), three level $\Delta$-learning (light blue), and multitask inference on $\Delta$s (gold). Reported performance is the average from six draws of each training set.
$\Delta$-learning is trained on differences between predictions of our quantity of interest while the single and multitask methods are trained on absolute predictions of the quantity of interest. Table 1 reports summary statistics for these data sets. In this case, the mean and variance of the differences between quantum chemistry methods are smaller by at least an order of magnitude than those of any set of quantum chemistry predictions. Since $\Delta$-learning is trained on differences to predict differences, it is expected that this method will produce smaller mean absolute error than a method trained on absolutes to predict absolutes.
By modifying our choice of tasks, we can design a multitask model which trains on and predicts differences. We call this the multitask $\Delta$ option. Fig. 5 compares this option to the multitask method on absolutes as well $\Delta$-learning. For each inference method, the average mean absolute error obtained for six random assignments of the training data is plotted for each of the same data set sizes which were considered in our previous experiments. The CAT secondary training set structure is used for the multitask absolute models as well as the multitask difference models. As expected, the multitask absolute model is not as successful as those which train on and learn differences.
In subplot (a) of Fig. 5, the object of estimation is the difference between CCSD(T) and PBE computations for each molecular system. For the multitask $\Delta$ option, we let the primary task be exactly this difference. The secondary task is the difference between SCAN and PBE computations which has correlation coefficient $0.782$ with the primary task data. Results of the multitask $\Delta$ option are plotted in gold. The basic $\Delta$-learning model is plotted in indigo and uses only two levels of theory, CCSD(T) and PBE, to predict their difference. Note that these models are the single task variant of the multitask $\Delta$ approach—they are trained only on the primary task. Therefore, Subplot (a) provides another example where the multitask method outperforms its single task counterpart, in this case with correlation parameter $\rho=0.782$. The turquoise points represent a $\Delta$-learning model trained with three levels of theory: CCSD(T), SCAN, and PBE. This model is the sum of a pair of two level models where the top level of the first difference—between SCAN and PBE—is the bottom level of the second—between CCSD(T) and SCAN. Given sufficiently large training sets for the lower levels of theory, this method is comparable with the multitask $\Delta$ option. The multitask $\Delta$ option outperforms most other models, particularly when either C or A is relatively small.
Subplot (b) of Fig. 5 shows the performance of the same set of methods when the primary and secondary tasks demonstrate a slight negative correlation. In this case, the extra data from the secondary task does not contribute to improved inference of the primary task. Here, we estimate the difference between CCSD(T) and SCAN predictions. For all descriptions accompanying Subplot (a), we reverse the roles of SCAN and PBE for Subplot (b). The major difference between the cases is the correlation coefficient between data for CCSD(T)-SCAN and PBE-SCAN: $\rho=-0.271$. The multitask models shown in Subplot (b) demonstrate the limitations of choosing Pearson’s coefficient as the correlation parameter. The multitask $\Delta$ option and three level $\Delta$-learning do not perform as well as two level $\Delta$-learning in this setting. This behavior suggests that the multitask method requires a stronger correlation between tasks to outperform the single task method. Further, the subplot shows that the multitask $\Delta$ option outperforms three level $\Delta$-learning for models trained on larger amounts of DFT data. Thus, three level $\Delta$-learning seems to suffer more from problematic data than the Multitask $\Delta$ option. Importantly, $\Delta$-learning has no correlation parameter and cannot adjust the contribution of each difference model. Though we have left a detailed optimization of the Multitask $\Delta$ option’s correlation parameter for future work, we observe that $\rho=0$ can always be chosen. Therefore, the fully optimized multitask $\Delta$ option will always perform at least as well as two level $\Delta$-learning. Consequently, the multitask framework makes it possible to benefit from useful secondary datasets which two level $\Delta$-learning cannot accommodate while offering more protection against uncorrelated data than three level $\Delta$ models.
In practice, we may have more than two tasks available to support inference, so we can further contrast the methods discussed in this section by considering how they scale to more tasks. Suppose that alongside PBE and CCSD(T) we want to include $n$ additional DFT functionals $y_{1},\dots,y_{n}$. Then, $\Delta$-learning requires us to put all quantum chemistry methods in some order—with CCSD(T) first and PBE last and the $y_{j}$ in between. The difference between neighboring pairs of methods is then learned. However, hierarchies in DFA accuracy can be difficult to predict and differ between problem settings. This necessary ordering can thus become rather arbitrary. By contrast, the multitask $\Delta$ option would incorporate the $n$ new secondary tasks based on the differences to PBE, i.e. $y_{j}-\text{PBE}$. Each of these secondary tasks can be directly related to the primary task through a correlation parameter—which thus eliminates the need to impose an arbitrary order.
IV.3 Ionization Potential of Small Organic Molecules
PBE0_DH PBE0 PBE BLYP
$\ \rho$ $0.970$ $0.967$ $0.954$ $0.955$
Table 2: Organic Molecules Case. Pearson’s correlation coefficients between each secondary task and the primary task data set. The column headers give the DFA used to generate the secondary set.
Fraction of shared molecules in CA: $0.0$
Refer to caption Refer to caption
Fraction of shared molecules in CA: $0.5$
Refer to caption Refer to caption
Fraction of shared molecules in CA: $1.0$
Refer to caption Refer to caption
(a) Trained without DFT data from Target Set (b) Trained with DFT data from Target Set
Figure 6: Organic Molecules Case. For different numbers of tasks (indicated by color), plots show MAE versus cost. Note that $10^{-0.3}\textrm{eV}$ is approximately half an electronvolt. For each row, a given fraction of all molecules in the C and A sets are used to generate training data for all secondary levels, and the remaining molecules are split among the secondary levels. The fractions are given at the top of each row. Models represented in column (a) were trained without DFT data from the T set while models in column (b) had access to this data.
[…]
In Fig. 6, as for previous results, the reported mean absolute error is the average over six tests produced by randomly assigning molecules to the C, A, and T sets. The left column of subplots compares results for multitask models trained on CA data sets to single task models trained only on CCSD(T) while the right column considers multitask models trained on CAT data. Translucent stripes encompass the results for models with different total numbers of tasks. For the two task case, the secondary task data is generated with PBE, and for the three task case the secondary data comes from PBE and PBE0. In all cases, secondary Target data is only generated with PBE. Recall that in our testing framework, when we train with the T data set, we train a custom model for each Target molecule. Specifically, the only T molecule in the training set is the molecule which we will use the model to predict. We remark that in practice it would not be necessary to train a new model for each molecule in the Target set.
As we observed in the water trimer case, the multitask models consistently perform better than single task models of comparable cost. Especially for relatively small training sets, the multitask method shows an improvement in performance when secondary Target data is included in the training set. Fig. 6 also informs us about the utility of including more than one secondary task in a model. Note that for each number of tasks we consider, we implement models with the same combinations of C, A, and T set sizes. In the top row, molecules in the C and A sets are each assigned to exactly one secondary task, regardless of the total number of secondary tasks. Therefore, in this row, for each model implemented with two total tasks, we consider models of the same cost implemented with three tasks and with five tasks. The results show that for sufficiently large training sets, there is a benefit to training with multiple tasks for a given budget. For smaller datasets, two tasks may lead to performance as good or better than a larger number of tasks. In particular, the top subplot in column (b) suggests that two task models perform particularly well when given a relatively inexpensive training set of the CAT structure. Note that, in each model, only one secondary task is given access to T data. Thus, in this setting, the decrease in performance when new tasks are added likely comes from siphoning off part of the computational budget to train tasks which do not have access to T data.
We can also consider whether it is useful to train on more than two predictions for the same molecular configuration. The middle row of Fig. 6 shows cases where each task has access to training data for the same 50% of molecules in the C set as well as 50% of the molecules in the A set while the remaining molecules are distributed one per task, and the final row shows cases where all tasks have access to all data in the C and A sets. From these subplots, we conclude that there is a benefit to training our model on multiple secondary predictions for the same molecule.
V Conclusion
The present work has demonstrated that inference models see a general benefit from leveraging all available data. Furthermore, use of the multitask framework can reduce the cost of generating entirely new training data by an order of magnitude and can also facilitate the opportunistic incorporation of existing datasets of heterogeneous quality into the training set. Our work is the first to explore a multitask GP regression model in a materials science setting. Significantly, the multitask framework is not restricted to the choice of GP inference, and future work can investigate training models with neural networks and other tools for regression.
Unlike other machine learning methods for constructing surrogates from ab initio data, the multitask model does not require all data sources be ordered in a hierarchy of accuracy. In this work, the data which informed our primary task—generated with CCSD(T)—was the most accurate in our training set, but we were also able to leverage an arbitrary number of secondary data sets without any information on their accuracy relative to each other. This feature of the multitask method is particularly useful when employing datasets constructed from different density functional approximations. We found that multitask model performance improves as we increase the number of secondary sets in the training data which include predictions for a molecule of interest. The influence of each secondary dataset is controlled by a correlation parameter, and we found that in most settings an estimate of Pearson’s correlation coefficient from training data works well. Use of this estimate simplifies implementation of the method. The multitask method also supports training on differences, and with appropriate choice of correlation parameter, multitask difference models can perform at least as well as and often better than $\Delta$ models. To explore the framework’s flexibility, we demonstrated the success of the multitask method on diverse training data set configurations.
Acknowledgements.
The authors would like to thank Yeongsu Cho, Chenru Duan, Dallas Foster, and Heather Kulik for insightful discussions. MIT SuperCloud and Lincoln Laboratory Supercomputing Center are acknowledged for providing the HPC resources that have contributed to the research results reported within this paper. Reuther et al. (2018) This material is based upon work supported by the Department of Energy, National Nuclear Security Administration PSAAP-III program, under Award Number DE-NA0003965 as well as work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 1745302. This research was supported by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602).
Author Contributions Statement
Katharine Fisher: Conceptualization (equal); Data Curation (equal); Methodology (equal); Software (lead); Writing – original draft (lead); Writing – review and editing (equal); Visualization (equal). Michael Herbst: Conceptualization (equal); Data Curation (equal); Methodology (equal); Supervision (equal); Writing – review and editing (equal); Visualization (equal). Youssef Marzouk: Conceptualization (equal); Methodology (equal); Supervision (equal); Writing – review and editing (equal); Visualization (equal).
Conflict of Interest Statement
The authors have no conflicts to disclose.
Data Availability Statement
All data supporting this work are openly available. All code to reproduce figures as well as all DFT predictions for the three-body energy of water trimers can be found at https://doi.org/10.5281/zenodo.10387647 or https://github.com/kefisher98/Multitask_GP/tree/Multitask. Water trimer configurations and CCSD(T) predictions for three-body energy are available at https://github.com/jmbowma/q-AQUA. The data that support the findings of this study for the small organic molecules example are at https://doi.org/10.5281/zenodo.10215421.
Appendix A Training Data Cost Model
In our numerical results, we estimate the training cost as
model training $\text{model training cost}\ =\ \sum_{j}^{m}n_{j}\hat{c}_{j}$
where $n_{j}$ is the number of training systems from the $j^{th}$ level of theory and $\hat{c}_{j}$ is an estimate of prediction cost for that level of theory. We produce one set of estimates for $\hat{c}_{1},\dots,\hat{c}_{m}$ for the water trimer example and a different set of estimates for the small organic molecules example.
For the water trimers example, we randomly selected ten molecular configurations from our data set as the basis for our cost model. Using Psi4, we computed the 3-b interaction energy at the level of SCAN and CCSD(T) for each configuration. We applied a counterpoise correction to ameliorate basis set superposition error. All computations for the water cost model where performed on physically the same Intel Xeon Platinum 8260 node of MIT SuperCloud (https://supercloud.mit.edu/systems-and-software). All computations for a given molecular configuration where performed on the same basis set subasis set sucore. For this case study, we estimate that a CCSD(T) calculation incurs $37$ times the cost of a DFT calculation based on the average prediction time of these methods on our example molecular configurations.
Several configurations in the small organic molecules data set are larger than the water trimer configurations. The median number of electrons in this data set is $40$, the minimum is $14$, and the maximum is $50$. Due to the expense of ionization potential calculations, we base our cost model on energy calculations for ten randomly sampled configurations with $36$ electrons. Note that CCSD(T) scales as $N^{7}$ where $N$ is the number of electrons in a system while DFT scales as $N^{3}$. Therefore, an estimate of the ratio between the cost of CCSD(T) and the cost of DFT based on $36$ electrons will be smaller than the actual ratio for most molecular configurations in our data set. Furthermore, the prediction of ionization potential was performed using ${\Delta}CC$ (delta coupled cluster). Thus, while our reported results (Fig. 6) estimate that CCSD(T) calculations demand $244$ times the expense of DFT, the true ratio between the cost of CCSD(T) and DFT is even larger.
Appendix B SOAP parameters
To fix reasonable values for parameters required by SOAP ($r_{cut}$, $\sigma_{atom}$, $n_{max}$, $l_{max}$), we refer to established conventions as well as experimentation. The cutoff radius, $r_{cut}$ controls the size of each local neighborhood, and a small value may lead to lost geometric insight. Unfortunately, surveys on feature construction Deringer et al. (2021); Musil et al. (2021) hold that a large cutoff radius does not necessarily provide proportionate insight: increasing radii larger than 6-8 Å is rarely if ever useful. Our choice of $\sigma_{atom}$ also influences our model of each atom’s neighborhood: it determines the lengthscale of the Gaussian shapes located on each of the surrounding atoms. The larger our choice of $\sigma_{atom}$, the more chance that Gaussian tails will slip passed the $r_{cut}$ border. Current literature Deringer et al. (2021) indicates that the best practice when working with the first three rows of the periodic table is to use $\sigma_{atom}=0.3\ $\mathrm{\SIUnitSymbolAngstrom}$$ for systems with hydrogen and $\sigma_{atom}=0.5\ $\mathrm{\SIUnitSymbolAngstrom}$$ for systems without. Finally, we choose the parameters $n_{max}$ and $l_{max}$ to control the size of our expansions of the local neighborhoods. Generally Deringer et al. (2021), choosing $n_{max}=12$ with $l_{max}=6$ is sufficient for high accuracy. In practice, the best values for $n_{max}$ and $l_{max}$ will be sensitive to the choice of radial basis set.
Figure 7: Example of inference results for a range of SOAP parameters. The colorbar gives the mean absolute error of predictions made for the ionization potential of small organic molecules. Both $r_{cut}$ and $\sigma_{atom}$ have units of $\mathrm{\SIUnitSymbolAngstrom}$.
We selected reasonable candidate values for each SOAP parameter and tested the performance of the resultant features in Gaussian process regression. All features for this work were calculated using the DScribe Python package. Himanen et al. (2020) Fig. 7 shows the mean absolute error (MAE) obtained with different SOAP parameters when GP regression trained with PBE level data is used to predict ionization potential (IP). Tests on CCSD(T) and PBE0 data produced similar results. Our tests include all combinations of $r_{cut}\in\{3,4,5\}$ , $\sigma_{atom}\in\{0.3,0.4,0.5\}$, $l_{max}\in\{2,4,6,8\}$, and $n_{max}\in\{6,8,10,12\}$. Units for $r_{cut}$ and $\sigma_{atom}$ are $\mathrm{\SIUnitSymbolAngstrom}$. We find that our accuracy is comparable to probabilistic predictions of IP found in literature. De et al. (2016) Of the SOAP parameters, the choice of $\sigma_{atom}$ has the largest impact on predictive performance.
For our numerical experiments in this work, we fix $\sigma_{atom}=0.4\ $\mathrm{\SIUnitSymbolAngstrom}$$ because this choice demonstrated the best overall performance across our tests. The other parameters are chosen to balance reasonable accuracy and cost. We fix both $l_{max}$ and $n_{max}$ at 8, and for the small organic molecules example, we use $r_{cut}=4\ $\mathrm{\SIUnitSymbolAngstrom}$$. For water trimers, we use $r_{cut}=10\ $\mathrm{\SIUnitSymbolAngstrom}$$ to capture all three component water molecules.
Appendix C Global Features
Figure 8: Correlation between different methods of constructing global features when calculating the distance between molecule pairs. De et al. (2016)
When using SOAP to compare two molecular systems, we must make modeling choices to standardize our representation of entire systems. SOAP features are constructed based on the neighborhood of each atom in a system, so challenges arise when two systems contain different numbers of constituent atoms of elements. There are multiple strategies for constructing “global” features to capture entire systems. De et al. (2016) The simplest approach is to average the SOAP features for each atom in the system. It is possible the loss of information from averaging may make it challenging to distinguish between the features of similar molecular systems. A variation to this approach averages the local features corresponding to each element contained in a molecular system. De et al. (2016) If we wish to compare systems A and B, and system B contains more elements than A, we insert SOAP features corresponding to isolated atoms of those excess elements in the representation of A. This model suggests that an atom of such an element does not interact with the rest of the system. A Regularized entropy match (REMatch) strategy has also been proposed for constructing global features. De et al. (2016) This approach solves a regularized optimization problem to find the match between the sets of local features for each molecular system which maximizes information entropy.
With the same set of small organic molecules which we use to test SOAP parameters, we investigate how well the global features constructed by different approaches agree. For pairs of molecules, we construct global features using each of the three approaches outlined in the previous paragraph. We represent difference between each pair by squaring an inner product of their features. This procedure is the same as using a polynomial kernel of degree 2. By comparing the output of the kernel for different featurization strategies, we determine whether the strategies generally agree about which molecules are similar.
Fig. 8 shows the correlation in kernel outputs corresponding to each pair of molecules for different global featurization strategies. Consistently, in our tests, we find strong correlation (Pearson’s coefficient >0.99)>0.99)> 0.99 ) between the REMatch features and features computed by averaging all local representations. These two strategies are also both positively correlated with the “Average by Species” approach which produces an averaged feature for each element in the system and inserts isolated atoms to represent elements not included in the system, but these correlations are not as strong as that between the REMatch and totally averaged features. We may see this result because the REMatch and totally averaged features share the same dimension, $n$, whereas features that are constructed by species specific averaging have dimension of $2n$ to $4n$ in the cases we considered. Because averaging features is much more computationally efficient than constructing REMatch features, we use this approach in our numerical experiments.

### v2

@@
 We demonstrate that multitask Gaussian process regression overcomes this limitation by leveraging both expensive and cheap data sources.
-In particular, we consider training sets constructed from coupled-cluster (CC) and density function theory (DFT) data.
-We report that multitask surrogates can predict at CC level accuracy with a reduction to data generation cost by over an order of magnitude.
+In particular, we consider training sets constructed from coupled-cluster (CC) and density functional theory (DFT) data.
+We report that multitask surrogates can predict at CC-level accuracy with a reduction to data generation cost by over an order of magnitude.
 Of note, our approach allows the training set to include DFT data generated by a heterogeneous mix of exchange-correlation functionals without imposing any artificial hierarchy on functional accuracy.
@@
 I Introduction
+model a data-based surrogate for quantum chemical prediction
+training data possibly heterogeneous data used to construct a model
+testing data data used to evaluate prediction error of a model; disjoint from training data
+level of theory the first principles method used to generate quantum chemical data (e.g., CCSD(T), DFT with PBE, …)
+high fidelity an accurate, expensive method such as CCSD(T)
+low fidelity a method, such as DFT, which is cheaper but less accurate than the high fidelity method
+task a model component trained with homogeneous data
+primary the task associated with the main prediction objective of the model
+secondary supplemental tasks which are included in a model to support prediction of the primary task
+core (C) molecular configurations for which training data is available for the primary task
+additional (A) molecular configurations for which training data is only available for secondary tasks
+target (T) molecular configurations for which the model will make predictions (i.e., the testing set)
+Table 1: Definitions of key terms.
 Predicting molecular properties based on first-principle approaches drives innovation across the physical sciences.
-However, exact prediction is infeasible due to the exponential scaling of the Schrödinger equation—the underlying quantum-mechanical description of materials systems.
-Decades of research have yielded numerous first-principle models to describe electronic structures approximately, each representing a compromise between computational cost and prediction accuracy.
+However, exact prediction is infeasible due to the exponential scaling of the Schrödinger equation—the underlying quantum-mechanical description of materials.
+Decades of research have yielded numerous first-principle methods to describe electronic structures approximately, each representing a compromise between computational cost and prediction accuracy.
 Coupled cluster (CC) methods are typical examples of high fidelity models which offer accuracy in exchange for considerable computational cost.
-Specifically, CCSD(T)—the “gold standard” of quantum-chemical prediction—suffers from a steep $O(N^{7})$ scaling where $N$ is the number of electrons in the target system.
+Specifically, CCSD(T)—the “gold standard” of quantum-chemical prediction—suffers from a steep $O(N_{e}^{7})$ scaling where $N_{e}$ is the number of electrons in the target system.
 Harding et al.
-(2008) In contrast, density-functional theory (DFT) methods are popular baseline models since they provide adequate accuracy for most systems while enjoying substantially lower cost (scaling only as $O(N^{3})$).
+(2008) In contrast, density functional theory (DFT) methods are popular baseline models since they provide adequate accuracy for most systems while enjoying substantially lower cost (scaling only as $O(N_{e}^{3})$).
 Ying, Yu, and Ying (2019); Perdew and Schmidt (2001); Goerigkab and Grimme (2011); Teale et al.
 (2022)
-Traditional workflows force users to settle on a single model which is cheap enough to result in manageable cost but still sufficiently high fidelity to give meaningful predictions.
-However, the advent of modern data-driven modeling has introduced new tools which make it possible to decouple expensive data generation from prediction.
-Statistical surrogates are trained with predictions from first-principle models and, once trained, can be substituted into workflows, making the prediction step considerably cheaper.
+Traditional workflows force users to settle on a single quantum chemical method that is cheap enough to result in manageable cost but still sufficiently high fidelity to give meaningful predictions.
+However, the advent of modern data-driven modeling has introduced new tools which make it possible to decouple the step of estimating the properties of unseen molecular systems from the step of running expensive first-principles computations.
+Statistical surrogate models are trained with predictions from first-principle methods and, once trained, can be substituted into workflows, making the prediction step considerably cheaper.
 In the context of building potential energy surfaces and interatomic potentials, these data-driven approaches have become the de facto standard.
@@
 (2020); Goodlett, Turney, and Schaefer (2023); Zaverkin et al.
-(2023) Simultaneously, large-scale data generation is facilitated by sophisticated workflow engines for high-throughput calculations Curtarolo et al.
+(2023) Simultaneously, large-scale data generation is facilitated by sophisticated workflows for high-throughput calculations Curtarolo et al.
 (2012); Jain et al.
@@
 (2022); Herbst and Levitt (2020, 2022); Cancès et al.
-(2023) The increasing amount of openly available data Ruddigkeit et al.
+(2023) Increasing amounts of data are openly available Ruddigkeit et al.
 (2012); Ramakrishnan et al.
 (2014); Chanussot* et al.
-(2021); Smith, Isayev, and Roitberg (2017) could in principle be readily employed, but quality is generally not homogeneous as each dataset’s author has made different modeling choices.
-Consequently, most surrogate models are single fidelity—all training data is generated using the same quantum-chemical method and the same numerical approach.
-Thus, heterogeneity has been an obstacle to opportunistically exploiting the already available data.
-With this work, we contribute to the promising development of multifidelity surrogates which do not require a single choice between accurate (but expensive) and fast (but crude) methods, but instead combine the advantages of both.
+(2021); Smith, Isayev, and Roitberg (2017) and are, in principle, ready for incorporation into statistical surrogates, but quality is generally not homogeneous as each dataset’s author has made different methodological choices.
+Consequently, most surrogate models are single fidelity: all training data is generated using the same quantum-chemical method and the same numerical approach.
+Heterogeneity has thus been an obstacle to opportunistically exploiting the already available data.
+In this work, we construct surrogates which do not require a single choice between accurate (but expensive) and fast (but crude) methods, but instead combine the advantages of both.
 Specifically, we consider multitask methods which construct a relationship between multiple regression problems without requiring precise information about the respective accuracy of each source.
 Bonilla, Chai, and Williams (2008); Leen, Peltonen, and Kaski (2012)
-Gaussian process (GP) regression is a popular tool for constructing surrogates informed by heterogeneous data.
+Figure 1: Diagram of training data sets compatible with the methods considered in this work.
+Here, we define three sets of molecules—core, additional, and target systems—and two fidelities—primary and secondary.
+Our goal is to train a model that will predict our quantity of interest for the target molecules with the accuracy of the primary method.
+The molecular space covered by secondary training data is a design choice for which we present six options, labeled C, CT, A, AT, CA, and CAT.
+Key terms are defined in Table 1.
+Our multitask framework can accommodate a range of heterogeneous training set structures.
+For example, suppose we aim to design a surrogate which predicts the three-body energy for any configuration of a water trimer with accuracy comparable to CCSD(T).
+We may only have access to $100$ CCSD(T) data points, all corresponding to configurations with HOH angles smaller $110$ degrees.
+Additional data may be provided by including a more exhaustive dataset based on DFT simulations with the PBE functional.
+Notably, even though the geometries of the molecular configurations in both datasets may not match, they can both be fully employed to train a multitask surrogate.
+Moreover, if higher predictive accuracy is desired, further training data can be created using yet another quantum-chemical method.
+For example, we may employ BLYP-based calculations on all molecular configurations covered by CCSD(T) and PBE, plus some configurations which we know we would eventually target for prediction.
+Remarkably, in this case, all available data can be used without imposing any assumption about which of the DFT methods (here, BLYP or PBE) is more accurate for the considered systems.
+Internally, a multitask surrogate creates a statistical submodel for each of the categories of data we supply it with (here, different levels of theory, i.e., CCSD(T), PBE, BLYP).
+In the language used in the rest of the article we call the submodel associated to our main prediction objective—the three-body energy at CCSD(T) level—the primary task.
+The other submodels—here associated to the DFT levels of theory— strive to improve the predictive accuracy of the primary task and are called secondary tasks.
+For a concise overview of these and other terms employed in the article, see Table 1.
+In this work, we will investigate such multitask models and the impact of training set design (i.e., the distribution of training data between the primary and secondary tasks) on prediction accuracy.
+Fig.
+1 visualizes the different training set structures which we consider.
+Secondary task training sets may share some molecular configurations with primary task training sets, include configurations unique to the secondary task, and may even include configurations from the testing data set.
+An example of the latter case is the use of a BLYP energy prediction at the same geometry where we later apply the surrogate to predict the CCSD(T) energy.
+The impact of these choices on inference is under-explored in literature, and multitask models are useful tools for investigating the benefits of different training set configurations.
+Crucially, the multitask framework enables training with a “dataset of opportunity,” amalgamated from existing data sets, as an alternative to expending computational time to generate entirely new data.
+In this paper, we construct submodels for each task using Gaussian process (GP) regression, a popular tool for constructing surrogates informed by heterogeneous data.
 Several works Pilania, Gubernatis, and Lookman (2017); Batra et al.
+(2019); Patra et al.
 (2020) have explored the use of multifidelity GP models Kennedy and O’Hagan (2000) trained using two fidelities, namely a high and a low fidelity DFT approach.
@@
 (2019); Dral, Zubatiuk, and Xue (2023).
-$\Delta$-learning can be extended to leverage more than two training data sets if multiple difference models are learned and summed.
+Hierarchical machine learning (hML) models extend the $\Delta$-learning framework to leverage more than two training data sets by learning and summing multiple difference models.
 Dral et al.
-(2020); Dral, Zubatiuk, and Xue (2023) However, between each pair of training sets, data points must be in some way aligned to allow for learning a difference model.
-For example, in a model of the difference between DFT and CCSD(T), DFT predictions may only contribute to the training set if a corresponding CCSD(T) prediction is also available.
-Further, when either $\Delta$-learning or multifidelity GP regression learn from more than two heterogeneous training sets, these sets must be assigned some order—for example, following a hierarchy in accuracy.
-Recent work Vinod, Kleinekathöfer, and Zaspel (2023) has proposed an optimized linear combination of submodels as an alternative, but all coefficients in this linear combination must be re-optimized whenever any submodel is altered, and opportunistic data set construction is not fully explored.
+(2020); Dral, Zubatiuk, and Xue (2023) However, between each pair of training sets, data points must be in some way aligned.
+For example, in a model of the difference between DFT and CCSD(T), DFT predictions may only contribute to the training set if the corresponding CCSD(T) prediction is also available.
+Further, when either hML or multifidelity GP regression learn from more than two heterogeneous training sets, these sets must be assigned some order—for example, following a hierarchy in accuracy.
+Recent work Vinod, Kleinekathöfer, and Zaspel (2023); Vinod et al.
+(2023) has proposed an optimized linear combination of submodels as an alternative, but all coefficients in this linear combination must be re-optimized whenever any submodel is altered, and opportunistic data set construction is not fully explored.
 Progress towards multifidelity approaches has also been made in the context of neural network surrogates.
-For example, transfer learning methods train an initial neural network on a larger source dataset, then re-optimise part of the parameters using a target dataset.
+For example, transfer learning methods train an initial neural network on a larger secondary dataset, then re-optimise part of the parameters using a primary dataset.
 Towell and Shavlik (1994); Fu (1989); Smith et al.
-(2019) For the prediction of atomistic properties, the source dataset might contain millions of DFT calculations while the target might be a smaller set of CCSD(T) predictions.
+(2019) For the prediction of atomistic properties, the secondary dataset might contain millions of DFT calculations while the primary might be a smaller set of CCSD(T) predictions.
 In prior studies, it has been observed that transfer learning can produce higher accuracy than training a network on only one of these datasets.
@@
 (2022), GP approaches are more amenable to mathematical analysis.
-Rasmussen and Williams (2006) This comparative transparency simplifies interpretation of the surrogate’s predictive process in the diverse training data configurations we analyze here.
+Rasmussen and Williams (2006) This comparative transparency simplifies interpretation of the surrogate’s predictive process in the diverse training data configurations we analyze.
 Additionally, GP methods have the advantage of yielding accurate predictions even for relatively small training set sizes.
-While the training cost of GP regression may become large as the size of the training set increases, many strategies for efficient scaling have been developed and can be routinely employed Quinonero-Candela and Rasmussen (2005); Liu et al.
-(2018); Wilson, Dann, and Nickisch (2015); Cole, Christianson, and Gramacy (2021).
-This work explores a GP regression approach which can accommodate an arbitrary number of quantum chemical datasets without an assumption of hierarchy.
-In particular, we employ the multitask framework Leen, Peltonen, and Kaski (2012), which has been shown to easily incorporate heterogeneous data sources.
-The notion of heterogeneity in fact encompasses many possibilities.
-Fig.
-1 compares the multitask method to other approaches by visualizing mappings between sets of molecules and the quantum chemistry methods which are used to generate viable training data sets.
-The highest level of theory used to construct these training sets will be called the primary level.
-In our case, the primary level will be CCSD(T).
-All other levels of theory will be categorized as secondary levels.
-Our secondary levels will include DFT methods using different density functional approximations (DFAs).
-Different levels of theory may contribute data points corresponding to the same set of molecules, different sets of molecules, or some union of shared and unique sets of molecules.
-Furthermore, the testing data set contains primary level predictions for target molecules.
-Thus, the multitask method has the option to use secondary predictions for the target molecules as part of the training set.
-The diagrams in Fig.
-1 show several options for constructing a training set for inference.
-The impact of these choices on inference is under-explored in literature, and the multitask method is a useful tool for investigating the benefits of different training set configurations.
-Crucially, the multitask framework enables training with a “dataset of opportunity,” amalgamated from existing data sets, as an alternative to expending computational time to generate entirely new data.
-Figure 1: Diagram of training data sets compatible with the methods considered in this work.
-Here, we define three sets of molecules—Core, Additional, and Target systems—and two quantum chemistry methods—primary and secondary.
-Our goal is to train a model that will predict our quantity of interest for the Target molecules with the accuracy of the primary method.
-By definition, the Core molecules are all systems for which we can compute training data with the primary method.
-Additional molecules are systems for which we have no primary data but which are not a part of the Target set.
-Theoretically, we may train using secondary predictions for molecules of all sets though not all inference methods support use of all data sets.
-The diagram shows the combinations of data sets and quantum chemistry methods supported by each inference method.
+While the training cost of GP regression may become large as the size of the training set increases, many strategies for efficient scaling have been developed and can be routinely employed.Quinonero-Candela and Rasmussen (2005); Liu et al.
+(2020); Wilson, Dann, and Nickisch (2015); Cole, Christianson, and Gramacy (2021)
 I.1 Advantages of the Multitask Method
@@
 We review the formulation of Gaussian process regression models.
-Subsections II.2 through II.4 discuss variants of GP regression which leverage heterogeneous sets.
+Subsections II.2 through II.5 discuss variants of GP regression which leverage heterogeneous data sets.
 See references Rasmussen and Williams (2006); Kennedy and O’Hagan (2000); Leen, Peltonen, and Kaski (2012); Bonilla, Chai, and Williams (2008); Abrahamsen (1997) for more detail.
 II.1 Gaussian Process Regression
-Our goal is to predict values of some quantity of interest (QoI) $f$, given an identifying input feature $\bm{X}$.
-We train our predictive model using $\{\bm{X}_{i},\bm{Y}_{i}\}_{i=1}^{n}$ where each $\bm{Y}_{i}$ is a noisy observation of $f(\bm{X}_{i})$.
+Our goal is to predict values of some quantity of interest (QoI) $f$, given an input feature vector $\bm{X}_{m}$ describing the $m^{th}$ molecular configuration.
+We train our predictive model using $\{(\bm{X}_{m},\bm{Y}_{m})\}_{{m}=1}^{M}$ where each $\bm{Y}_{m}$ is a noisy observation of $f(\bm{X}_{m})$.
 In our setting, we interpret imperfect first-principle predictions of the QoI as noisy observations of the true quantity.
-Each $\bm{X}_{i}$ is some representation of the electronic structure of a molecular system.
-For our numerical results, we build $\bm{X}_{i}$ using the well established Smooth Overlap of Atomic Potentials (SOAP).
+Each $\bm{X}_{m}$ is some representation of the electronic structure of a molecular system.
+For our numerical results, we build $\bm{X}_{m}$ using the well established Smooth Overlap of Atomic Potentials (SOAP).
 Bartók et al.
-(2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) $\bm{Y}_{i}$ is the CCSD(T) or DFT prediction of the quantity of interest for the $i^{th}$ molecular configuration.
+(2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) $\bm{Y}_{m}$ is the CCSD(T) or DFT prediction of the quantity of interest for the ${m}^{th}$ molecular configuration.
 We will assume that each observation is independently perturbed by additive noise which follows identical Gaussian distributions. The corresponding model is
-$\bm{Y}_{i}$ $=$ $\ f(\bm{X}_{i})\ +\ \varepsilon_{i},$ (1)
-$\varepsilon_{i}$ $\sim$ $\ N\left(0,\ \sigma^{2}\right).$
-Before inference, we choose a mean function, $\mu(\bm{X})$ to describe the expected value of $f(\bm{X})$ and a kernel function, $k(\cdot,\cdot)$, to encapsulate our assumptions about the relationship between $f(X_{i})$ and $f(X_{j})$.
+$\bm{Y}_{m}$ $=$ $\ f(\bm{X}_{m})\ +\ \varepsilon_{m},\quad m=1,\dots,M,$ (1)
+$\varepsilon_{m}$ $\sim$ $\ N\left(0,\ \sigma^{2}\right).$
+Before inference, we choose a mean function $\mu(\bm{X}_{m})$ to describe the expected value of $f(\bm{X}_{m})$ and a kernel function, $k(\cdot,\cdot)$, to encapsulate our assumptions about the relationship between $f(\bm{X}_{m})$ and $f(\bm{X}_{m^{\prime}})$.
 Typically, $k(\cdot,\cdot)$ will have hyperparameters that are selected through optimization, then fixed during inference.
@@
 $f(\bm{X})\ \sim\ \text{GP}\Big{(}\mu(\bm{X}),k(\bm{X},\bm{X}^{\prime})\Big{)}.$ (2)
-As a consequence of this prior assumption, any finite collection, $\{f(\bm{X}_{1}),\dots,f(\bm{X}_{n}\})$, will follow a multivariate normal distribution with mean $\bm{\mu}=\{\mu(\bm{X}_{1}),\dots,\mu(\bm{X}_{n})\}$ and covariance matrix $K$ with entries at row $i$ and column $j$ given by $k(\bm{X}_{i},\bm{X}_{j})$.
-We are interested in inferring the values the function takes on at target inputs, and we collect these values in the vector, $\bm{f}_{*}$.
-Using this notation, we write the joint distribution on the observations and targets for inference:
+As a consequence of this prior assumption, any finite collection, $\{f(\bm{X}_{1}),\dots,f(\bm{X}_{M})\}$, will follow a multivariate normal distribution with mean $\bm{\mu}=\{\mu(\bm{X}_{1}),\dots,\mu(\bm{X}_{M})\}$ and covariance matrix $K$ with entries at row $m$ and column $m^{\prime}$ given by $k(\bm{X}_{m},\bm{X}_{m^{\prime}})$.
+It is common practice to center the observation vector and set $\bm{\mu}=\bm{0}$.
+We retain a more general formulation here to emphasize that it is possible to construct a feature-dependent prior mean.
+We are interested in inferring the values the QoI takes on at target inputs, and we collect these values in the vector, $\bm{f}_{*}$.
+Using this notation, we write the joint distribution of the observations and targets for inference:
 $\begin{bmatrix}\bm{Y}\\ \bm{f}_{*}\end{bmatrix}\ \sim\ N\left(\begin{bmatrix}\bm{\mu}\\ \bm{\mu}_{*}\end{bmatrix},\ \begin{bmatrix}K_{pp}+\sigma^{2}I&K_{p*}\\[3.0pt] K_{p*}^{T}&K_{**}\end{bmatrix}\right).$ (3)
 where $I$ is the identity matrix.
-The specific subscripts of each covariance matrix represent the features that the matrix compares: $p$ corresponds to our primary training data and *** indicates the target set.
-We note that in this single-fidelity setting all training data is contained in the primary set, but this will no longer be the case when we consider surrogates informed by heterogeneous data.
+The specific subscripts of each covariance matrix represent the features that the matrix compares: $p$ corresponds to our primary training data and * indicates the target set.
+We note that in this single task setting, all training data is contained in the primary set, but this will no longer be the case when we consider surrogates informed by heterogeneous data.
 The marginal distribution of $\bm{Y}$ is the model evidence, and conditioning $\bm{Y}$ on $\bm{f}_{*}$ yields the likelihood.
@@
 The predicted covariance indicates how much our observations have reduced our uncertainty, provided our assumptions are reasonably aligned to reality.
-II.2 $\Delta$-Learning
-Suppose that for each training input, we have data on our QoI from multiple sources.
-We thus seek an inference model that leverages $\left\{\bm{Y}_{ij}\right\}_{i=1,j=1}^{n,m}$, where we have data for molecular configurations $i=1,\dots,n$ and methods $j=1,\dots,m$.
-A simple “pooling” approach may put all observations into one vector and proceed with inference via Eq.
+II.2 Multiple Observation Sets
+Suppose we have access to heterogeneous training data which includes predictions of our QoI by multiple first principles methods.
+We thus seek an inference model that leverages all data $\bigcup_{n=1}^{N}\{\bm{Y}_{mn}\}_{m\in\mathcal{M}_{n}}$, where each $n=1,\dots,N$ represents a task and the corresponding set $\mathcal{M}_{n}$ contains indices of the molecular configurations available to train a model for task $n$.
+A simple “pooling” approach could put all observations into one vector and proceed with inference via Eq.
 (4).
-Then, $\bm{Y}_{ij_{1}}$ and $\bm{Y}_{ij_{2}}$ are interpreted as independent observations of the QoI, $f(\bm{X}_{i})$, for molecular configuration $\bm{X}_{i}$.
-The distinct behavior of each quantum chemical method is treated as a perturbation of some mean value by a realization of the noise distribution.
-We seek instead a model where the distribution of each $\bm{Y}_{ij}$ is sensitive to $j$.
-Ideally, we want to map differences in electronic structures methods to differences in their predictions.
+Then, if $m\in\mathcal{M}_{n}\cap\mathcal{M}_{n^{\prime}}$, we can interpret $\bm{Y}_{mn}$ and $\bm{Y}_{mn^{\prime}}$ as independent observations of the same QoI, $f(\bm{X}_{m})$, for molecular configuration $\bm{X}_{m}$.
+The distinct behavior of each quantum chemical method is treated as a perturbation of some mean value by a realization of the noise.
+Rather than doing this, we seek a model where the distribution of each $\bm{Y}_{mn}$ is sensitive to $n$.
+Ideally, we want to map differences in electronic structure methods to differences in their predictions.
 Method-specific features could be a solution, but construction of such features is challenging.
@@
 This strategy puts a large burden on covariance hyperparameter optimization since one set of hyperparameters must encode differences between molecular configurations and between quantum chemical methods.
-Alternatively, to characterize the qualitative difference between methods within the features, we would have to solve the very problem for which we are creating the features in the first place—for instance, we would need to reliably predict CCSD(T) performance given CCSD(T) and DFT data from training molecules as well as DFT data for target molecules.
-Instead, we use this subsection to discuss $\Delta$-learning, which directly models the differences between the predictions of quantum chemistry methods.
-$\Delta$-learning performs Gaussian process regression to predict differences, $\Delta_{i}$, between observations obtained with two different methods for a given $\bm{X}_{i}$.
+Alternatively, to characterize the qualitative difference between methods within the features, we would have to solve the very problem for which we are creating the features in the first place: for instance, we would need to reliably predict CCSD(T) performance given CCSD(T) and DFT data from training molecules as well as DFT data for target molecules.
+Instead, we will extend the framework of GP regression to model relationships between multiple prediction tasks.
+II.3 $\Delta$-Learning
+$\Delta$-learning performs Gaussian process regression to predict differences, $\Delta_{m}$, between observations obtained with two levels of theory for a given $\bm{X}_{m}$.
+In this setting, the observation data available to us allows two interpretations: either as a single-task approach with the observations $\{(\bm{Y}_{m2}-\bm{Y}_{m1})\}_{m=1}^{M}$, or as a two-task approach with data $\bigcup_{n=1}^{2}\{Y_{mn}\}_{m\in\mathcal{M}_{n}}$.
+Unless stated otherwise, we will assume the latter.
 If we retain our assumption that noise follows an independent, identically distributed Gaussian distribution, we have the model
-$\bm{Y}_{ij_{1}}-\bm{Y}_{ij_{2}}$$\equiv$ $\ \Delta_{i}\ =\ f\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{i},$ (11)
-$f$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}_{\Delta},\ k_{\Delta}(\bm{X},\bm{X}^{\prime})\big{)}.$
-where $j_{1}$ and $j_{2}$ indicate two prediction methods.
+[…]
+$=$ $\ f_{\Delta}\big{(}\bm{X}_{m}\big{)}\ +\ \varepsilon_{m}$
+$f_{\Delta}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{\Delta}(\bm{X}),\ k^{\Delta}(\bm{X},\bm{X}^{\prime})\big{)}.$
+where $n$ and $n^{\prime}$ indicate two prediction methods.
 An application of Eq.
-(10) yields a distribution on the regression function for $\Delta$.
+(4) yields a distribution on the regression function for $\Delta(n,n^{\prime})$.
 Suppose our ultimate interest is to predict $\bm{Y}$ at $\bm{X}_{*}$.
-Within the $\Delta$ framework, we can supply an observation from one method, say $\bm{Y}_{*j_{2}}$, in order to make a prediction for the value from the other method:
-$\bm{Y}_{*j_{1}}\ \approx\ \Delta_{i}\ +\ \bm{Y}_{*j_{2}}.$
-Thus, $\Delta$-learning is suited to settings where calculations by method $j_{1}$ are both more accurate and more time consuming than calculations that use $j_{2}$.
-Consider a case where we want to build a model from an existing data set which includes predictions from CCSD(T) and several DFT methods.
-If the predictions all correspond to the same set of molecules, $\Delta$-learning models have the ability to leverage more of the data set than conventional GP regression models.
+Within the $\Delta$-learning framework, we can supply an observation from one method, say $\bm{Y}_{*n^{\prime}}$, to make a prediction for the value from the other method:
+$\bm{Y}_{*n}\ \approx\ \Delta_{*}\ +\ \bm{Y}_{*n^{\prime}}.$
+Thus, $\Delta$-learning is suited to settings where calculations by method $n$ are both more accurate and more time consuming than calculations that use $n^{\prime}$.
+Consider a case where we want to build a model from an existing data set which consists of CCSD(T) predictions for configurations in $\mathcal{M}_{2}$ and DFT predictions for configurations in $\mathcal{M}_{1}$.
+In the notation introduced by Table 1, $\mathcal{M}_{2}$ is the core set, $C$.
+$\mathcal{M}_{1}$ may contain any combination of molecules from the $C$, $A$, and $T$ sets.
+As (5) indicates, the training data for $\Delta$-learning will include only molecular configurations in $\mathcal{M}_{1}\cap\mathcal{M}_{2}\subseteq C$.
+If $\mathcal{M}_{1}\cap\mathcal{M}_{2}=C$, then $\Delta$-learning will be able to leverage more data than the conventional GP regression model described by (3) trained on the CCSD(T) data.
 Unfortunately, $\Delta$-learning is not-applicable in cases where predictions by different methods correspond to different molecular configurations.
-To use a DFT prediction for a molecule $\bm{X}_{i}$ to train a $\Delta$ model between CCSD(T) and DFT, we also require a CCSD(T) prediction for the same molecule.
+When $\mathcal{M}_{1}\cap\mathcal{M}_{2}=\emptyset$, it is impossible to train a $\Delta$-learning model.
+To include a DFT prediction for a molecule $\bm{X}_{m}$ in the training set for a $\Delta$ model between CCSD(T) and DFT, we also require a CCSD(T) prediction for the same molecule.
+Hierarchical machine learning (hML)Dral et al.
+(2020); Dral, Zubatiuk, and Xue (2023) extends the $\Delta$-learning approach to $N>2$ by representing the difference between $\bm{Y}_{mN}$ and $\bm{Y}_{m1}$ as
+$\Delta^{hML}_{m}(N,1)\ =\ \sum_{n=1}^{N-1}\Delta_{m}(n+1,n),$ (6)
+where each $\Delta_{m}(n+1,n)$ is defined following (5) (with $f_{\Delta}$ potentially different for each $n$).
+As before, to ultimately predict, $\bm{Y}_{*N}$, we also require data for $\bm{Y}_{*1}$.
+Note that the summed difference models are independently trained, and their training sets need not be identical.
+Thus, an hML model designed to predict at the level of CCSD(T) would not necessarily require CCSD(T) data for every molecular configuration, $m$, in its training set.
+Rather, each training configuration must belong to some intersection of two adjacent levels of theory, i.e., to $\bigcup_{n=1}^{N-1}\left(\mathcal{M}_{n}\cap\mathcal{M}_{n+1}\right)$.
 In the following subsections, we consider alternative approaches which model disparities in data sources and have the ability to use all available observations.
-II.3 Multifidelity Method
-As a first step toward considering statistical models which train on heterogeneous data, we describe the multifidelity data fusion model.
+II.4 Multifidelity Method
+As a step toward considering statistical models which train on heterogeneous data, we describe the multifidelity data fusion model.
 Kennedy and O’Hagan (2000) The multifidelity approach defines two regression functions: $f_{p}$, representing the predictions from some high fidelity model, and $f_{s}$, representing a low fidelity model.
 Kennedy and O’Hagan (2000) We assume the relationships
-$\bm{Y}_{ip}$ $=$ $\ f_{p}\big{(}\bm{X}_{i}\big{)}\ +\varepsilon_{ip}$ (12)
-$=$ $\ \rho f_{s}\big{(}\bm{X}_{i}\big{)}\ +\ \delta_{ps}\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{ip},$
-$\bm{Y}_{is}$ $=$ $\ f_{s}\big{(}\bm{X}_{i}\big{)}\ +\ \varepsilon_{is}.$ (13)
-The term $\delta_{ps}$ captures the difference between the high fidelity model and the low fidelity model, scaled by hyperparameter, $\rho$.
-This disparity is endowed with a Gaussian process prior, as is the low fidelity regression function, $f_{s}$.
+$\bm{Y}_{mp}$ $=$ $\ f_{p}\big{(}\bm{X}_{m}\big{)}\ +\varepsilon_{mp},\qquad\forall\,m\in\mathcal{M}_{p}$ (7)
+$=$ $\ \rho f_{s}\big{(}\bm{X}_{m}\big{)}\ +\ \delta_{ps}\big{(}\bm{X}_{m}\big{)}\ +\ \varepsilon_{mp},$
+$\bm{Y}_{ms}$ $=$ $\ f_{s}\big{(}\bm{X}_{m}\big{)}\ +\ \varepsilon_{ms},\qquad\forall\,m\in\mathcal{M}_{s}.$ (8)
+The term $\delta_{ps}$ captures the difference between the high fidelity model and the low fidelity model, scaled by the hyperparameter $\rho$.
+This disparity is endowed with a Gaussian process prior, as is the low fidelity regression function, $f_{s}$:
+$\delta_{ps}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{\delta}(\bm{X}),\ k^{\delta}(\bm{X},\bm{X}^{\prime})\big{)},$
+$f_{s}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{s}((\bm{X})),\ k^{s}(\bm{X},\bm{X}^{\prime})\big{)}.$
 The correlation, $\rho$, will be optimized before inference.
-As before, we assume that the additive noise terms, $\varepsilon_{ip}$ and $\varepsilon_{is}$, are independent and drawn from identical, centered Gaussian distributions.
-Now, suppose we want to predict our targets, $\bm{f}_{*}$, at the high fidelity level.
-We can determine that the joint distribution of our high and low fidelity observations and our targets has covariance
+As before, we assume that the additive noise terms, $\varepsilon_{mp}$ and $\varepsilon_{ms}$, are independent and drawn from identical, centered Gaussian distributions.
+Now, suppose we want to predict our QoI for the target molecules, $\{\bm{X}_{m}:m\in T\}$, at the primary fidelity.
+Let $\bm{f}_{*}$ be the vector of all values in $\{f_{p}(\bm{X}_{m}):m\in T\}$.
+We can determine that the joint distribution of our high and low fidelity observations and $\bm{f}_{*}$ is multivariate normal with mean and covariance
+$\mu$ $=$ $\ \begin{bmatrix}\bm{\mu}^{s}\ +\ \bm{\mu}^{\delta}\\ \bm{\mu}^{s}\\ \bm{\mu}^{s}\ +\ \bm{\mu}^{\delta}\\ \end{bmatrix}$
 $\Sigma$ $=$ $\ \begin{bmatrix}\rho^{2}K^{s}_{pp}+K^{\delta}_{pp}+\sigma^{2}I&\rho K^{s}_{ps}&\rho^{2}K^{s}_{p*}+K^{\delta}_{p*}\\[5.0pt] \rho K^{s}_{sp}&K^{s}_{ss}+\sigma^{2}I&\rho K^{s}_{s*}\\[5.0pt] \rho^{2}K^{s}_{*p}+K^{\delta}_{*p}&\rho K^{s}_{*s}&\rho^{2}K^{s}_{**}+K^{\delta}_{**}\\[5.0pt] \end{bmatrix}.$ (9)
 where $\sigma^{2}$ is the variance of the noise distribution.
-The superscript of each covariance matrix, K, indicates which Gaussian process prior corresponds to the kernel function used in its construction, and the subscripts indicate the pairs of features that are compared.
+The superscript of each mean vector $\bm{\mu}$ and covariance matrix, K, corresponds to the GP prior used in its construction, and the subscripts indicate the pairs of features that are compared by the prior kernel function.
 Bayesian inference yields an analytical posterior distribution, analogous to Eq.
 (4).
-II.4 Multitask Method
-“Multitasking” refers to a category of methods which consider several Gaussian process regression tasks and assume some relationship between these tasks.
+II.5 Multitask Method
+“Multitasking” refers to a category of methods which consider several regression tasks and assume some relationship between these tasks.
 As with $\Delta$-learning or multifidelity fusion, multitasking can allow us to incorporate a larger dataset into our inference problem without just pooling the data into one observation vector.
-For instance, we can define a regression problem trained on CCSD(T) data as well as one trained on each DFA we consider.
+For instance, we can define a regression problem trained on CCSD(T) data as well as one trained on each density-functional approximations (DFA) we consider.
 By modeling correlation between regression functions, we can use the DFA regression tasks to support prediction in the CCSD(T) task.
-We can apply multitask regression to problems with $m$ tasks where, for each task $j\in\{1,\ldots,m\}$, we have observational data $\bm{Y}_{j}\in\mathbb{R}^{n}$.
-In practice, the method can be generalized, so that observation sets of different sizes are used for different tasks.
 Symmetric variants of multitask modeling treat all tasks equally.
-For instance, we may assume that all the data are modeled by a single Gaussian process whose covariance kernel is the Kronecker product of one kernel which relates the tasks and another which models relationships between data from the same task.Bonilla, Chai, and Williams (2008)
-In many cases, however, the goal is to perform regression for a primary task, and all other tasks may be considered secondary tasks which provide data that may be useful in learning the primary task.
+For instance, we may assume that all the data are modeled by a single Gaussian process whose covariance kernel is the Kronecker product of one kernel which relates the tasks and another which models relationships between data from the same task.Bonilla, Chai, and Williams (2008); Khatamsaz, Vela, and Arróyave (2023) In many cases, however, the goal is to perform regression for a primary task, and all other tasks are considered secondary tasks which provide data that may be useful in learning the primary task.
 For our own setting, CCSD(T) is a natural choice of data for the primary regression task because we would like to make predictions which match this method in accuracy.
@@
 Thus, we consider an asymmetrical model Leen, Peltonen, and Kaski (2012) in which secondary tasks are related through the primary task.
-Suppose $\bm{Y}_{p}\in\mathbb{R}^{n_{p}}$ is the observation data for the primary task and $\bm{Y}_{s_{j}}\in\mathbb{R}^{n_{j}}$ is the data for the $s_{j}^{th}$ secondary task.
+Let $\mathcal{M}_{p}$ contain the molecular configurations available to train the primary task.
+Note that following the definitions provided in Table 1, $\mathcal{M}_{p}=C$.
+Likewise, we define $\mathcal{M}_{s_{n}}$ to be the set of training configurations for the $n^{th}$ secondary task.
 As before, we assume that each task has its own regression function
-$\bm{Y}_{ip}$ $=$ $\ f_{p}(\bm{X}_{i})\ +\ \varepsilon_{ip},$ (18)
-$\bm{Y}_{is_{j}}$ $=$ $\ f_{s_{j}}(\bm{X}_{i})\ +\ \varepsilon_{is_{j}}\qquad\forall j=1,\dots,m.$ (19)
+$\bm{Y}_{mp}\ =\ \ f_{p}(\bm{X}_{m})\ +\ \varepsilon_{mp},$ $\forall\,m\,\in\mathcal{M}_{p},$ (10)
+$\bm{Y}_{ms_{n}}\ =\ f_{s_{n}}(\bm{X}_{m})\ +\ \varepsilon_{ms_{n}},$ $\forall\,m\,\in\mathcal{M}_{s_{n}},$
+$\forall\,n=1,\dots,N-1.$
 where each noise term is drawn independently from a centered Gaussian distribution with variance $\sigma^{2}$.
 We now model a shared structure in the regression functions based on the primary function:
-$f_{s_{j}}(\bm{X}_{i})$ $=\ \rho_{s_{j}}f_{p}(\bm{X}_{i})\ +\ \delta_{s_{j}}(\bm{X}_{i})\qquad\forall j=1,\dots,m.$ (20)
-The specific component, $\delta_{s_{j}}$, of the secondary function serves to “explain away” behavior that is not captured by the shared component, $f_{p}$een, Peltonen, and Kaski (2012) Both the correlation parameter, $\rho_{s_{j}}$ and the specific component aim to mitigate negative transfer—learning behavior from secondary tasks that is not representative of the primary task.
-We make the prior assumption that $f_{p}$ and $\delta_{s_{j}}$ for $j=1,\dots,m$ follow Gaussian process distributions, each with a mean and kernel function which we will specify.
-As was the case in subsections II.1 through II.3, the first step toward inference with the multitask method is to determine the joint distribution on observations and targets.
-Note that the specific component of each task is modeled as independent of the others.
-Then, the covariance between observations due to the specific component is a block matrix with zeros corresponding to the covariance of specific functions for different tasks.
-Call this block matrix $K^{\delta}$.
-Let $\bm{R}$ be a diagonal matrix with dimensions equal to the number of observations, and let each diagonal element be the correlation hyperparameter of the corresponding task.
-The covariance of the joint distribution is
+$f_{s_{n}}(\bm{X}_{m})$ $=\ \rho_{s_{n}}f_{p}(\bm{X}_{m})\ +\ \delta_{s_{n}}(\bm{X}_{m}).$ (12)
+The task-specific component $\delta_{s_{n}}$ of the secondary function serves to “explain away” behavior that is not captured by the shared component, $f_{p}$een, Peltonen, and Kaski (2012) Both the correlation parameter $\rho_{s_{n}}$ and the task-specific component aim to mitigate negative transfer—learning behavior from secondary tasks that is not representative of the primary task.
+We make the prior assumption that $f_{p}$ and $\delta_{s_{n}}$ for $n=1,\dots,N-1$ follow Gaussian process distributions, each with a mean and kernel function which we will specify:
+$f_{p}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{p}(\bm{X}),\ k^{p}(\bm{X},\bm{X}^{\prime})\big{)},$ (13)
+$\delta_{s_{n}}$ $\sim$ $\ \text{GP}\big{(}\bm{\mu}^{s_{n}}(\bm{X}),\ k^{s_{n}}(\bm{X},\bm{X}^{\prime})\big{)}.$ (14)
+As was the case in subsections II.1 through II.4, the first step toward inference of $\bm{f}_{*}$ with the multitask method is to determine the joint distribution of observations and targets.
+Since all secondary tasks are modeled following (12), we can easily represent this distribution for arbitrary $N$.
+Note that the task-specific component $\delta_{s_{n}}$ of each secondary task is modeled as independent of the others, $\delta_{s_{n^{\prime}}}$ for $n\neq n^{\prime}$.
+Then, the covariance between observations due to the tasks-specific components is a block diagonal matrix, which we denote as $K^{\delta}$.
+The kernel function $k^{s_{n}}$ is applied to construct the $s_{n}^{th}$ diagonal block.
+Since the off-diagonal blocks correspond to the covariances of task-specific specific functions for different tasks, they are all zero.
+We note that the number of rows and columns of this matrix both equal $\sum_{n=1}^{N-1}|\mathcal{M}_{s_{n}}|$.
+Let $\bm{R}$ be a diagonal matrix with dimensions equal to the number of secondary observations, and let each diagonal element be the correlation hyperparameter of the corresponding task.
+The mean and covariance of the multivariate normal joint distribution is
+$\mu$ $=$ $\ \begin{bmatrix}\bm{\mu}^{p}\\ R\bm{\mu}^{p}+\bm{\mu}^{s}\\ \bm{\mu}^{p}\\ \end{bmatrix}$
 $\Sigma$ $=$ $\ \begin{bmatrix}K^{p}_{pp}+\sigma^{2}I&K^{p}_{ps}\bm{R}&K^{p}_{p*}\\[5.0pt] \bm{R}K^{p}_{sp}&\bm{R}K^{p}_{ss}\bm{R}+K^{\delta}+\sigma^{2}I&\bm{R}K^{p}_{s*}\\[5.0pt] K^{p}_{*p}&K^{p}_{*s}\bm{R}&K^{p}_{**}\\[5.0pt] \end{bmatrix}.$ (15)
-The superscript $p$ indicates that a matrix was created with the primary kernel, and the matrix subscripts indicate the features that are compared. Inference proceeds as described in Eq. (10).
-II.5 Beyond Two Tasks
+Superscripts indicate the GP prior structures used to create a vector or matrix, and subscripts indicate the features that are compared by a matrix.
+In both cases, $s$ indicates evaluations for all secondary tasks, $n=1,\dots,N-1$.
+Inference proceeds as described in Eq.
+(4).
+II.6 Comparison of Methods for More than Two Fidelities
 Asymmetric multitasking assumes a relationship between regression functions—given by Eq.
@@
 $\vdots$
-$f_{s_{j-1}}(\bm{X})$ $=$ $\ \rho_{s_{j}}f_{s_{j}}(\bm{X})\ +\ \delta_{s_{j}}(\bm{X}),$
+$f_{s_{N-1}}(\bm{X})$ $=$ $\ \rho_{s_{N}}f_{s_{N}}(\bm{X})\ +\ \delta_{s_{N}}(\bm{X}),$
 whereas in the asymmetric approach, all functions share the same relationship to the primary function
@@
 $\vdots$
-$f_{s_{j}}(\bm{X})$ $=$ $\ \rho_{s_{j}}f_{p}(\bm{X})\ +\ \delta_{s_{j}}(\bm{X}).$
+$f_{s_{N}}(\bm{X})$ $=$ $\ \rho_{s_{N}}f_{p}(\bm{X})\ +\ \delta_{s_{N}}(\bm{X}).$
 Thus, compared to multifidelity fusion, the asymmetric multitasking approach more naturally accommodates a reasonably large number of regression tasks and relates each closely to the primary task of interest.
-Multitasking also holds a similar advantage over $\Delta$-learning.
-Before the notion of a difference can be imposed it is necessary to choose some order between the source data sets.
+Multitasking also holds a similar advantage over the hML extension to $\Delta$-learning.
+The indices $n=1,\dots,N$ suggest an order of the training sets, and the difference models in (6) correspond to consecutive pairs of training sets.
 Such an order may be arbitrary, for example, if our data sets are generated by multiple DFAs with similar structure.
 In Section III, we will describe different compositions of a training data set which the multitask framework allows us to test.
-III Heterogeneous Training Data
+III Mapping Heterogeneous Data to Tasks
+$\mathcal{C}(t)\quad t\in\{p,s_{1},\dots,s_{N-1}\}$ intersection between core molecules and training molecules for task $t$
+$\mathcal{A}(t)\quad t\in\{p,s_{1},\dots,s_{N-1}\}$ intersection between additional molecules and training molecules for task $t$
+$\mathcal{T}(t)\quad t\in\{p,s_{1},\dots,s_{N-1}\}$ intersection between target molecules and training molecules for task $t$
+Table 2: Definition of functions which return sets of molecular configurations corresponding to primary task, p, and secondary tasks, $s_{1},\dots,s_{N-1}$.
+Note that by definition, $\mathcal{C}(p)$ is the entire core set, $C$, and $\mathcal{A}(p)=\mathcal{T}(p)=\emptyset$.
 In our numerical examples, we train inference models using data for multiple molecular configurations computed with multiple quantum chemistry methods.
 The multitask method leaves many questions about the particulars of the training data set up to the researcher.
-For instance, available data must be apportioned to different tasks.
-One of our goals is to demonstrate that by training with data from multiple quantum chemistry methods, we can construct an inference model with prediction accuracy comparable to CCSD(T) with less computational expense than if we trained a single task model on CCSD(T) alone.
-Thus, our CCSD(T) data is the primary target of our regression problem and a natural choice to inform the primary task.
-We let each secondary task correspond to a different DFA.
-Examples with one up to four secondary tasks will be considered.
-Another decision to be made is whether the training data sets for different tasks should include the same molecular configurations.
-In other words, we must determine whether or not to train with predictions for the same quantity by different methods.
-If our target for inference is the CCSD(T) prediction of energy for a system of water molecules, we may also choose to include DFT predictions of this energy in the training set for a secondary task.
-The relative benefits of constructing training sets with different degrees of intersection between tasks is under-explored in literature.
-Our work seeks to understand the impact of training set design on model performance.
-We consider a range of possible training set compositions here, and Fig.
-2 provides two example diagrams of our system for distinguishing between different constructions.
-We assign the molecular configurations to three sub data sets—Core (C), Additional (A), and Target (T).
-Dots in the figure indicate when configurations are available at a given level of theory.
-Black indicates availability to the training set, and green indicates membership of the testing set.
-The C set contains all configurations for which CCSD(T) training data is available, the A set contains configurations for which we have DFT data but no CCSD(T) data, and the T set contains configurations that we use to test our model’s ability to reproduce CCSD(T) level accuracy.
+In particular, it is necessary to define tasks and map molecules to the training set of each task.
+III.1 Selecting Tasks
+The primary task is chosen according to our prediction goal.
+We aim to demonstrate that by training with data from multiple quantum chemistry methods, we can construct an inference model which has prediction accuracy comparable to CCSD(T) for less computational expense than if we trained a single task model on CCSD(T) alone.
+Thus, we will test our surrogates using a set of CCSD(T) predictions for some target molecules belonging to a set we call $T$.
+Our primary task must be informed by CCSD(T) predictions for a different set of molecules which we refer to as our core training set, $C$.
+See Table 1 for an overview of our key definitions.
+We consider two approaches to mapping between levels of theory and tasks.
+In the first (subsections IV.2.1 and IV.3), the training data of our primary task will consist of direct computations of our QoI at CCSD(T) level, and each secondary task will be trained by DFT predictions of the QoI obtained with different DFT functionals.
+The second approach (Subsection IV.2.2) we call the “multitask $\Delta$” option.
+In that setting, the levels of theory include CCSD(T) and DFT computations performed with two different DFAs.
+The difference between CCSD(T) and the first DFA will be the primary task while the difference between the two DFAs informs the secondary task.
+Note that the $\Delta$-learning method could be interpreted as the single task counterpart to the multitask $\Delta$ option.
+III.2 Assigning Molecular Configurations to Tasks
+We are specifically interested in the flexibility which the multitask framework provides towards assigning molecular configurations to secondary training sets.
+Our pool of available molecules is divided into Core ($C$), Additional ($A$), and Target ($T$) subsets to systematically explore the impact of the molecular space covered by the secondary training data and its relationship to our primary training and testing data.
+The $A$ set contains all molecular configurations that are not part of the primary training or testing set.
+In our setting, these are configurations which are not our targets for inference and for which high level CCSD(T) predictions are unavailable.
+A major difference between the multitask framework and $\Delta$-learning is the former’s ability to leverage this $A$ set.
+Fig.
+1 shows the six main secondary training set structures considered in this work.
+We name the training set structures using combinations of the letters $C$, $A$, and $T$.
+Inclusion of a letter in a structure name indicates that at least one molecular configuration from the corresponding subset is included in the secondary training data set.
+For example, a CT training structure in our setting would contain CCSD(T) predictions for the C set of molecules, DFT predictions for some subset of the C set, and DFT predictions for some subset of the T set.
+The data requirement of implementing a multitask method with this CT training set structure is comparable to the data demands of a $\Delta$-learning model.
+However, a crucial distinction is that the multitask approach does not require a DFT prediction for every molecular configuration in the C and T sets.
+Furthermore, the multitask model can make use of an additional $A$ training set for which CCSD(T) predictions are unavailable.
+$\Delta$-learning models cannot be implemented for such a training set structure.
+III.3 Multiple Secondary Tasks
+The multitask framework can accommodate an arbitrary number of secondary tasks and allows the observational data used to train each task to correspond to non-identical sets of inputs.
+We will use the same definition for the $C$, $A$, and $T$ molecular sets regardless of the total number of tasks, $N$.
+In this context, when a training structure is labeled CAT, this name indicates that at least one configuration from each of these sets is present in the training set for at least one secondary task.
+To the best of our knowledge, most work which explores training set design for multifidelity, $\Delta$-learning, or hML models focuses on optimizing the size of the training set for each source.
+Questions of whether different secondary sources of information should explore the same or new areas of molecular space are underexplored in literature.
+For example, if our secondary training data includes DFT predictions at the PBE level for CO2 configurations and at the BLYP level for ONH4 configurations, we aim to determine how much our model will improve if we fill in the blanks (i.e., add PBE level predictions for ONH4 and BLYP level predictions for CO2 versus adding data on previously unseen molecules).
+Such questions are investigated in Subsection IV.3.
 Figure 2: Example data set structures.
@@
 The vertical alignment of dots indicates that multiple tasks have access to the same molecular configurations.
-In our numerical results, we always train models with the CCSD(T) data from the C set and consider various compositions of the secondary sets.
-All secondary tasks are highlighted in gray.
-We identify different training set compositions by the letters of the sets of molecule which are included in our secondary training data.
-For example, the top diagram in Fig.
-2 shows an A training set because the only DFT predictions available to train the model come from the Additional set.
-By contrast, the lower diagram shows a CAT training set because dots are present in the slots for DFT predictions for Core, Additional, and Target molecular configurations.
-Furthermore, the vertical alignment of dots indicate when some of the same molecular configurations are present in the training sets for multiple tasks.
-In the CAT example in Fig.
-2, the dots indicate that the secondary tasks share $20\%$ of the C and A sets while the rest of the molecules in these sets are assigned to only one task each.
-In Section IV, we will show results for the C, CT, A, AT, CA, and CAT training set configurations.
-Fig.
-6 shows the impact of varying the fraction of molecules that the secondary tasks have in common.
-Our testing framework is designed so that models trained with DFT Target data can readily be compared with equivalent models that do not train on this data.
-Since the inference model can make Target predictions one at a time, for each Target molecule we train a new model using the C set, A set, and the DFT prediction for only that Target molecule.
-While in practice we could take the much more efficient approach of training on the entire DFT Target set at once, our approach enables us to contrast each CAT model to a CA model with a training set which only differs by one molecule.
-Consequently, we can isolate the effect of the DFT level Target data on prediction performance.
-In Section IV, we will examine the performance of the $\Delta$ and Multitask methods, informed by CCSD(T) and DFT predictions, and compare them to basic Gaussian process regression, informed only by CCSD(T).
+To make this discussion more tractable, we introduce notation which allows us to compare the sets of molecules used to train each task.
+Table 2 defines three set-valued functions: $\mathcal{C}$, $\mathcal{A}$, and $\mathcal{T}$.
+To illustrate this notation further, Figure 2 provides a visual representation of two possible training set structures.
+Each line in the figure corresponds to a different task, here defined by the level of theory employed during data generation.
+CCSD(T) informs the primary task, and different DFT functionals correspond to each secondary task (shaded in grey).
+Black dots represent training data, and their location indicates whether molecules from a given set are available for a particular task.
+Vertically aligned dots indicate the presence of the same molecular configuration in the training set for multiple tasks.
+For example, consider the top plot of Figure 2: SCAN and PBE are the secondary tasks, $s_{1}$ and $s_{2}$, and we have:
+$\bigl{|}\mathcal{C}(s_{n})\bigr{|}$ $=$ $\ 0,\quad n\in{1,2}$
+|$\frac{\bigl{|}\{\mathcal{A}(s_{1})\cap\mathcal{A}(s_{2})\}\bigr{|}}{|A|}$ $=$ $\ 1,$
+where ||$\cdot$|| ⋅ | returns the number of elements in a given set, and $A$ is the entire ‘additional’ set.
+The training data for both secondary tasks only includes molecules from the A set, so this plot represents an A test case.
+The lower plot represents a CAT data structure where  $\forall\,1\leq n<n^{\prime}\leq 4$ and
+|$\frac{\bigl{|}\{\mathcal{C}(s_{n})\cap\mathcal{C}(s_{n^{\prime}})\}\bigr{|}}{|C|}\ =\ \frac{\bigl{|}\{\mathcal{A}(s_{n})\cap\mathcal{A}(s_{n^{\prime}})\}\bigr{|}}{|A|}\ =\ \frac{2}{10}.$
+III.4 Training with Target Molecules
+We stress that whenever $\mathcal{T}(s_{n})\neq\emptyset$, the secondary training data includes low level predictions of at least some molecules which we will use to test the multitask model.
+In Fig.
+2, the testing molecules—target molecules at the primary fidelity—are represented in green.
+The black dots in the third column of the lower plot indicate that BLYP predictions for the test molecules are included in the training set.
+All lower level predictions for target molecular configurations which are available at training time can be used to obtain a multitask model.
+If new lower level T predictions become available after this initial training, a retraining is needed to use this data.
+Note that a $\Delta$-learning model can leverage such additional T data without any retraining.
+However, each target data point will only contribute to a single prediction of the $\Delta$-learning framework.
+In contrast, retraining the multitask model buys the benefit of this additional data point for all future predictions.
+Furthermore, sparse update techniques such as the Sherman-Morrison formula offer efficient alternatives to a full retraining of the multitask model.
 IV Numerical Examples
@@
 IV.1.1 Data Generation
-For our first test case—inference of 3-b interaction energy—5986 water trimer configurations and corresponding CCSD(T) calculations were randomly selected from 45,332 configurations available from https://github.com/jmbowma/q-AQUAYu et al.
-(2022) Density functional theory data was generated for 2992 of these configurations using Psi4.
-Smith et al.
-(2020) We employed the PBEPerdew, Burke, and Ernzerhof (1996) and SCANSun, Ruzsinszky, and Perdew (2015) functional approximations.
-The former was selected for its popularity in materials science research and the latter because its description of intermediate range dispersion has been found to contribute to accurate prediction of water energy differences Chen et al.
+Multitask models can be trained on heterogeneous data sets constructed opportunistically from multiple sources without the need to specify a predefined accuracy hierarchy (apart from singling out the primary task).
+Our test cases were selected to explore this flexibility by tapping mostly on previously published data sets supplemented with only few additional computations.
+For our first test case—inference of 3-b interaction energy—we consider 5986 water trimer configurations and corresponding CCSD(T) calculations.
+These were randomly selected from 45,332 configurations available from https://github.com/jmbowma/q-AQUA Yu et al.
+(2022) to obtain training data sizes approximately matching our second test case.
+To illustrate that even such truncated datasets can be employed to yield reasonable predictive qualities with multitask models, we enrich this data set again by running a small number of cheap density-functional theory simulations: on 2992 of these configurations we employ Psi4 Smith et al.
+(2020) to compute the 3-b interaction energies at PBEPerdew, Burke, and Ernzerhof (1996) and SCANSun, Ruzsinszky, and Perdew (2015) level.
+The former was selected for its popularity and the latter because its description of intermediate range dispersion has been found to contribute to accurate prediction of water energy differences Chen et al.
 (2017); Dasgupta et al.
@@
 We base our cost model for each method on the average runtime of Psi4 computations for ten representative systems.
-For more details about the cost model, see Appendix A.
+For more details about the cost model, see Appendix B.
+We make the code required to reproduce our results and figures openly available as described in Statement Data Availability Statement.
+Similarly, the first principles data as well as surrogate prediction data are openly accessible along with the cost of training and applying surrogate models.
 IV.1.2 Features
@@
 (2021); Musil et al.
-(2021) More details are available in Appendices B and C.
+(2021) More details are available in Appendices C and D.
 IV.1.3 Hyperparameter Optimization
-Squared exponential kernels are used within all Gaussian process priors in our experiments.
-Abrahamsen (1997) These kernels introduce mean, variance ($v$), and length scale ($\ell$) hyperparameters in addition to the multitask correlation ($\rho$).
+Squared exponential kernels are used within all Gaussian process priors in our experiments: Abrahamsen (1997)
+$k(\bm{X},\bm{X}^{\prime})\ =\ v\ \exp\left(\frac{-(\bm{X}-\bm{X}^{\prime})^{T}(\bm{X}-\bm{X}^{\prime})}{2\ell}\right)$
+These kernels introduce mean, variance ($v$), and length scale ($\ell$) hyperparameters in addition to the multitask correlation ($\rho$).
 We use the mean of our training data for each task as that task’s prior mean, calculated at training time.
@@
 Sóbester (2007) For the organic molecules example, we also use a maximum likelihood estimate for $v$ while for the water trimer example, we find better results from using an estimate of $v$ which minimizes mean absolute prediction error.
+See Appendix A for a comparison of MLE and minimum MAE estimates as well as a more detailed discussion of the resources required for hyperparameter optimization.
+IV.1.4 Isolating the Impact of Target Training Data
+Consider a molecular configuration, $\bm{X}_{m}$.
+We are interested in contrasting the accuracy of a multitask model trained using a CA set and a model trained on a CAT set which is identical except for the inclusion of a DFT prediction for $\bm{X}_{m}$.
+The Target set may contain a large number of molecular configurations, and we cannot make a fair comparison if we simultaneously include DFT predictions for all Target molecules in the training set.
+To facilitate comparison, whenever $|\mathcal{T}(s_{n})|>0$ for $n=1,\dots,N-1$ we train a new multitask model for each element of $\mathcal{T}(s_{n})$.
+Each model will train on only one configuration from the $T$ set, and we will test that model by making predictions for the same configuration.
+We report the mean absolute error (MAE) over all these models and compare this value to the MAE of the predictions of a CA trained model for the same set of targets.
+We implement this retraining procedure to investigate the impact of training set design.
+Retraining is not a prerequisite for leveraging all molecules in $\mathcal{T}(s_{n})$.
+All lower level data for target molecules which is available at training time can be simultaneously incorporated into one multitask model.
+See Subsection III.4 for further discussion.
 IV.2 Three-Body Interaction Energy of Water Trimer
@@
 The legends of the plots indicate the structure of the secondary training set according to the system presented by Fig.
-2.
-All MAE reported are the average results obtained from six randomly chosen training sets.
-We first compare the performance of multitask models with only one secondary task to the performance of a traditional GP—or single task—model.
+1.
+The red arrows emphasizes the order of magnitude improvement in data generation cost achieved by CAT multitask models, quantified in Table 3.
+We first compare the performance of multitask models with only one secondary task ($N=2$) to the performance of a traditional GP model ($N=1$) .
 Fig.
 3 shows the error incurred by several implementations of each model when predicting the three-body (3-b) energy of water trimer configurations.
-We used the SCAN functional to generate the secondary task data, and Pearson’s correlation coefficient between the primary and secondary tasks is $0.997$, indicating strong positive correlation.
+We used the SCAN functional to generate the secondary task data.
+Pearson’s correlation coefficient between the primary and secondary task training data is $0.997$, indicating strong positive correlation.
 The three subplots of Fig.
@@
 All models were tested on a target set of 320 molecular configurations.
-We let the size of the Core set, $N_{c}$, take on values in $\{5,10,20,40,80,160,320\}$ for both inference methods and additionally consider $N_{c}\in\{640,1280,2560\}$ for the single task method.
-The subplots of Fig.
-3 also show seven clear groupings of data points with steep slope.
-This effect is a consequence of the high cost of CCSD(T) relative to DFT.
-Each of the seven groupings corresponds to a different $N_{c}$.
-Given the size of the Core set, we let the number of configurations in the A set range from $0.5N_{c}$ to $6N_{c}$.
-Each point on the plot represents a model with a different training set size.
-The x axis indicates the cost of the training set while the y axis gives the average mean absolute error for six random assignments of molecular configurations to the C, A, and T sets.
-We provide details of our cost model in Appendix A.
+For multitask models, we consider seven possible sizes of the core set: $|C|\in\{5,10,20,40,80,160,320\}$.
+We use the same seven core sets as well as sets with $|C|\in\{640,1280,2560\}$ to train the corresponding single task ($N=1$) models.
+Performance of this single task reference is marked with a black line on each subplot.
+Larger $|C|$ is considered in the single task case to demonstrate the cost required for a these models to reach the accuracy of the best performing multitask models.
+We choose the additional training sets for multitask models so that $|A|=r|C|$ where $r$ is allowed to take on each value in $\{0.5,1,2,3,4,5,6\}$.
+Whenever core data is included in the secondary training set, we only consider $\mathcal{C}(s_{1})=C$.
+Thus, we consider a total of $7$ different training set sizes for the C and CT structures and $49$ different training set sizes for the A, AT, CA, and CAT structures.
+Each point on each subplot of Fig.
+3 represents a model with a different training set size.
+The x axis indicates the cost of each model’s training set; details of our cost model are provided in Appendix B.
+The y axis gives the average value of mean absolute error obtained from six models trained on a given training set size and structure.
+For each of these six implementations, available molecular configurations are randomly assigned to the C, A, and T sets without replacement.
+For every secondary training set structure except C, the multitask method outperforms the single task reference.
+Including T data from DFT in the training set leads to better results than corresponding cases without T in every example.
+In particular, training with secondary T data can enable the multitask method to exhibit an accuracy comparable to a GP model that is an order of magnitude more expensive.
+Table 3 provides representative examples.
+The data points corresponding to the last line in the table are marked with arrows on Fig.
+3.
+This analysis of MAE demonstrates that each statistical model can achieve chemical accuracy.
+For insight into prediction for individual molecules, we report that the best performing multitask models produce predictions with $\geq 95\%$ correlation to the CCSD(T) target data.
+Accuracy [meV] Cost [millions of seconds]
+MAE Single Task CA CAT
+$14$ $2.0$ $0.58$ $0.13$
+$7$ $16.0$ $4.5$ $2.0$
+Table 3: Water Trimer. Data generation cost to achieve the threshold of accuracy given in the left column.
 Refer to caption Refer to caption
@@
 Figure 4: Water Trimer Case.
-Effect of including Core and Additional data in secondary training sets.
+Effect of including Core and Additional data (defined in Table 1) in secondary training sets.
 Each indigo point represents a comparison between two multitask models where one model uses a training set which is a subset of the other.
@@
 The $y=x$ line is plotted in green for reference.
-For every training set configuration except C, the multitask method outperforms the single task reference, shown by a black line.
-Including T data from DFT in the training set leads to better results than corresponding cases without T in every example.
-In particular, training with secondary T data can enable the multitask method to exhibit an accuracy comparable to a GP model that is an order of magnitude more expensive.
 While the C results demonstrate no significant improvement on the single task model, both A and CA results show steady improvement as the size of the A set increases.
-This behavior suggests that for the multitask method to be beneficial, secondary training data must cover molecular space that is not included in the CCSD(T) training data.
+We can distinguish the impact of increasing $|A|$ from the impact of increasing $|C|$ in Fig.
+3 because CCSD(T) is significantly more expensive than DFT.
+Thus, every jump in $|C|$ leads to a significant increase in cost while jumps in $|A|$ for fixed $|C|$ produce a comparatively slight increase to cost.
+The positive correlation between $|A|$ and performance whenever $\mathcal{T}(s_{1})=\emptyset$ suggests that for the multitask method to be beneficial, secondary training data must cover molecular space that is not included in the CCSD(T) training data.
 We mention that it is possible that the behavior of C models would improve if we consider task dependent features or different relationships between task regression functions.
@@
 Furthermore, unlike the AT model, the CAT model does not show improvement as the size of the A set increases.
-It seems that for our setting, the combination of secondary C and T training data is so valuable to the model that any benefit from A data is dwarfed.
-We can hypothesize that when the model has access to data for the Core molecules from the primary and secondary tasks, it implicitly learns information about the difference between the tasks.
-Given secondary data on the Target, it can leverage the captured relationship between tasks to make an accurate prediction for the Target on the primary level.
+It seems that for our setting, the combination of secondary C and T training data is so valuable to the model that any benefit from the A set is dwarfed.
+We can hypothesize that when the model has access to training data for the Core molecules from both the primary and secondary tasks, it implicitly learns information about the difference between the tasks.
+Given secondary data on the Target, it can leverage the captured relationship between tasks to make an accurate prediction for the Target on the primary fidelity.
 Fig.
@@
 While Fig.
-3 demonstrates that multitask models can be successful when trained on many different training data configurations from the CAT framework, it does not directly compare the performance of the A and CA models.
+3 demonstrates that multitask models outperform single task models for several training set configurations, it does not directly compare the efficacy of the A and CA training structures.
 Such a comparison will illuminate the resilience of the multitask model to the loss of points in the training set which are common to primary and secondary tasks.
@@
 Subplot (b) of Fig.
-4 also compares models trained with and without C secondary data in the case where all of these models also train on secondary Target data.
+4 also compares models trained with and without C secondary data in the case where $\mathcal{T}(s_{1})\neq\emptyset$ for all models.
 While the AT models perform well, there is a penalty to losing access to secondary C training data when secondary Target data is also available.
@@
 absolute median $0.0126$ $0.0144$ $0.0155$
-variance $0.0009$ $0.0010$ $0.0011$
+standard deviation $0.0300$ $0.0316$ $0.0332$
 CCSD(T)--SCAN [eV] CCSD(T)--PBE [eV]
@@
 absolute median $0.0013$ $0.0018$
-variance $1.1\times 10^{-5}$ $3.2\times 10^{-5}$
-Table 1: Water Trimer case. Summary statistics for absolute quantum chemistry predictions and for differences.
+standard deviation $0.0033$ $0.0057$
+Table 4: Water Trimer case. Summary statistics for training data.
 (a) Estimation of CCSD(T)-PBE
@@
 Table 4 reports summary statistics for these data sets.
-In this case, the mean and variance of the differences between quantum chemistry methods are smaller by at least an order of magnitude than those of any set of quantum chemistry predictions.
-Since $\Delta$-learning is trained on differences to predict differences, it is expected that this method will produce smaller mean absolute error than a method trained on absolutes to predict absolutes.
+In this case, both the mean and standard deviation of the differences between quantum chemistry methods are smaller by at least an order of magnitude than than the corresponding statistics for any set of quantum chemistry predictions.
+Greater standard deviation in the distribution of target data tends to result in greater standard deviation in the prediction error of a statistical surrogate which in turn leads to larger MAE.
+Since $\Delta$-learning is trained on differences to predict differences, it is thus expected that this method will produce smaller mean absolute error than a method trained on absolutes to predict absolutes.
 By modifying our choice of tasks, we can design a multitask model which trains on and predicts differences.
-We call this the multitask $\Delta$ option.
+This reinterpretation of tasks will be referred to as the multitask $\Delta$ option.
 Fig.
-5 compares this option to the multitask method on absolutes as well $\Delta$-learning.
-For each inference method, the average mean absolute error obtained for six random assignments of the training data is plotted for each of the same data set sizes which were considered in our previous experiments.
-The CAT secondary training set structure is used for the multitask absolute models as well as the multitask difference models.
+5 compares this option to $\Delta$-learning as well as our original implementation of the multitask model, trained on absolute predictions of the QoI.
+For each inference method, we report the average MAE obtained over six random assignments of the training data to the C, A, and T sets.
+We consider the same molecule set sizes which were considered in our previous experiments.
+The CAT secondary training set structure is used for the multitask absolute QoI models as well as the multitask $\Delta$ option models.
 As expected, the multitask absolute model is not as successful as those which train on and learn differences.
@@
 Results of the multitask $\Delta$ option are plotted in gold.
-The basic $\Delta$-learning model is plotted in indigo and uses only two levels of theory, CCSD(T) and PBE, to predict their difference.
-Note that these models are the single task variant of the multitask $\Delta$ approach—they are trained only on the primary task.
-Therefore, Subplot (a) provides another example where the multitask method outperforms its single task counterpart, in this case with correlation parameter $\rho=0.782$.
-The turquoise points represent a $\Delta$-learning model trained with three levels of theory: CCSD(T), SCAN, and PBE.
-This model is the sum of a pair of two level models where the top level of the first difference—between SCAN and PBE—is the bottom level of the second—between CCSD(T) and SCAN.
-Given sufficiently large training sets for the lower levels of theory, this method is comparable with the multitask $\Delta$ option.
-The multitask $\Delta$ option outperforms most other models, particularly when either C or A is relatively small.
+$\Delta$-learning predictions for CCSD(T)-PBE are plotted in indigo.
+Note that these $\Delta$-learning models are the single task counterpart to the multitask $\Delta$ approach—they are a conventional GP model trained only on its primary task.
+Therefore, Subplot (a) provides another example where the multitask method outperforms its single task counterpart, in this case with a smaller correlation parameter than we considered in Subsection IV.2.1: $\rho=0.782$.
+The turquoise points represent a hierarchical machine learning (hML) extension to the $\Delta$-learning model, trained with three levels of theory: CCSD(T), SCAN, and PBE.
+In this method, the difference between CCSD(T) and PBE is represented as
+[…]
+$\coloneqq$ $\ \Delta(\text{CCSD(T)},\ \text{SCAN})\ +\ \Delta(\text{SCAN},\ \text{PBE}),$
+where the submodels $\Delta$(CCSD(T), SCAN) and $\Delta$(SCAN, PBE) are $\Delta$-learning models as defined in (5).
+Note that while $\Delta$(CCSD(T), SCAN) can only be trained with molecular configurations in $C$, $\Delta$(SCAN, PBE) can be trained with $C\cup A$.
+Given sufficiently large training sets for SCAN and PBE, this method is comparable with the multitask $\Delta$ option.
+The multitask $\Delta$ option outperforms most other models, particularly when either $|C|$ or $|A|$ is relatively small.
 Subplot (b) of Fig.
-5 shows the performance of the same set of methods when the primary and secondary tasks demonstrate a slight negative correlation.
+5 shows the performance of the same set of methods when the primary and secondary tasks do not have significant correlation.
 In this case, the extra data from the secondary task does not contribute to improved inference of the primary task.
@@
 The multitask models shown in Subplot (b) demonstrate the limitations of choosing Pearson’s coefficient as the correlation parameter.
-The multitask $\Delta$ option and three level $\Delta$-learning do not perform as well as two level $\Delta$-learning in this setting.
-This behavior suggests that the multitask method requires a stronger correlation between tasks to outperform the single task method.
-Further, the subplot shows that the multitask $\Delta$ option outperforms three level $\Delta$-learning for models trained on larger amounts of DFT data.
-Thus, three level $\Delta$-learning seems to suffer more from problematic data than the Multitask $\Delta$ option.
-Importantly, $\Delta$-learning has no correlation parameter and cannot adjust the contribution of each difference model.
+The multitask $\Delta$ option and hML extension to $\Delta$-learning do not perform as well as standard $\Delta$-learning in this setting.
+Since standard $\Delta$-learning is the single task counterpart to the multitask $\Delta$ option, this behavior suggests that the multitask method requires a stronger correlation between tasks to outperform the single task method.
+Further, the subplot shows that the multitask $\Delta$ option outperforms hML for models trained on larger amounts of DFT data.
+Thus, hML seems to suffer more from problematic data than the Multitask $\Delta$ option.
+Importantly, $\Delta$-learning and its hML extension have no correlation parameters and cannot adjust the contribution of each level of theory.
 Though we have left a detailed optimization of the Multitask $\Delta$ option’s correlation parameter for future work, we observe that $\rho=0$ can always be chosen.
-Therefore, the fully optimized multitask $\Delta$ option will always perform at least as well as two level $\Delta$-learning.
-Consequently, the multitask framework makes it possible to benefit from useful secondary datasets which two level $\Delta$-learning cannot accommodate while offering more protection against uncorrelated data than three level $\Delta$ models.
-In practice, we may have more than two tasks available to support inference, so we can further contrast the methods discussed in this section by considering how they scale to more tasks.
-Suppose that alongside PBE and CCSD(T) we want to include $n$ additional DFT functionals $y_{1},\dots,y_{n}$.
-Then, $\Delta$-learning requires us to put all quantum chemistry methods in some order—with CCSD(T) first and PBE last and the $y_{j}$ in between.
+Therefore, the fully optimized multitask $\Delta$ option will always perform at least as well as standard $\Delta$-learning.
+Consequently, the multitask framework makes it possible to benefit from useful secondary datasets which standard $\Delta$-learning cannot accommodate while offering more protection against uncorrelated data than hML models.
+We conclude our comparison between hML and the multitask $\Delta$ option with a remark about scaling when training data is available for more levels of theory.
+Suppose that alongside PBE and CCSD(T) we want to include $L$ additional DFT functionals $y_{1},\dots,y_{L}$.
+Then, an hML model which uses all levels of theory would require that these levels be assigned some order—with CCSD(T) first and PBE last and $y_{1},\dots,y_{L}$ in between.
 The difference between neighboring pairs of methods is then learned.
@@
 This necessary ordering can thus become rather arbitrary.
-By contrast, the multitask $\Delta$ option would incorporate the $n$ new secondary tasks based on the differences to PBE, i.e.
-$y_{j}-\text{PBE}$.
+By contrast, the multitask $\Delta$ option would incorporate the $L$ new secondary tasks based on the differences to PBE, i.e., $y_{\ell}-\text{PBE}$.
 Each of these secondary tasks can be directly related to the primary task through a correlation parameter—which thus eliminates the need to impose an arbitrary order.
@@
 Table 5: Organic Molecules Case. Pearson’s correlation coefficients between each secondary task and the primary task data set. The column headers give the DFA used to generate the secondary set.
-Fraction of shared molecules in CA: $0.0$
+$k=0$
 Refer to caption Refer to caption
-Fraction of shared molecules in CA: $0.5$
+$k=0.5$
 Refer to caption Refer to caption
-Fraction of shared molecules in CA: $1.0$
+$k=1$
 Refer to caption Refer to caption
@@
 Note that $10^{-0.3}\textrm{eV}$ is approximately half an electronvolt.
-For each row, a given fraction of all molecules in the C and A sets are used to generate training data for all secondary levels, and the remaining molecules are split among the secondary levels.
-The fractions are given at the top of each row.
+For each row of plots, $k$ determines the fraction of training data which is shared between any pair of secondary sets.
 Models represented in column (a) were trained without DFT data from the T set while models in column (b) had access to this data.
 […]
+|$\frac{\bigl{|}\{\mathcal{C}(s_{n})\cap\mathcal{C}(s_{n^{\prime}})\}\bigr{|}}{|C|}\ =\ \frac{\bigl{|}\{\mathcal{A}(s_{n})\cap\mathcal{A}(s_{n^{\prime}})\}\bigr{|}}{|A|}\ =\ k.$
+We provide results for $k\in\{0,\frac{1}{2},1\}$. These cases correspond—in order, from the top down—to the three rows of Fig. 6.
 In Fig.
-6, as for previous results, the reported mean absolute error is the average over six tests produced by randomly assigning molecules to the C, A, and T sets.
-The left column of subplots compares results for multitask models trained on CA data sets to single task models trained only on CCSD(T) while the right column considers multitask models trained on CAT data.
-Translucent stripes encompass the results for models with different total numbers of tasks.
+6, as for previous results, the reported mean absolute error is the average result of six tests produced by randomly assigning molecules to the C, A, and T sets.
+The left column of subplots compares results for multitask models trained using the CA structure to the single task reference while the right column considers multitask models trained with the CAT structure.
+Translucent stripes encompass the results for models with a given number of tasks, $N$.
 For the two task case, the secondary task data is generated with PBE, and for the three task case the secondary data comes from PBE and PBE0.
 In all cases, secondary Target data is only generated with PBE.
-Recall that in our testing framework, when we train with the T data set, we train a custom model for each Target molecule.
-Specifically, the only T molecule in the training set is the molecule which we will use the model to predict.
+Recall from Subsection IV.1.4 that our training and testing procedure is designed to produce fair comparisons between models trained on CA and CAT training structures.
+Though we test the predictive performance of each training set structure on $320$ different target molecular configurations, we do not simultaneously include all of these configurations in $\mathcal{T}(\text{PBE})$.
+Rather, we iterate over the target configurations and train a multitask model for each configuration such that $\mathcal{T}(\text{PBE})=1$ in each case.
 We remark that in practice it would not be necessary to train a new model for each molecule in the Target set.
@@
 Especially for relatively small training sets, the multitask method shows an improvement in performance when secondary Target data is included in the training set.
+Further, we report that for larger secondary training sets, correlation between multitask predictions and the CCSD(T) target data exceeds $95\%$, but we defer a full evaluation of prediction quality at the level of individual molecular configurations to future work.
 Fig.
 6 also informs us about the utility of including more than one secondary task in a model.
-Note that for each number of tasks we consider, we implement models with the same combinations of C, A, and T set sizes.
-In the top row, molecules in the C and A sets are each assigned to exactly one secondary task, regardless of the total number of secondary tasks.
-Therefore, in this row, for each model implemented with two total tasks, we consider models of the same cost implemented with three tasks and with five tasks.
-The results show that for sufficiently large training sets, there is a benefit to training with multiple tasks for a given budget.
-For smaller datasets, two tasks may lead to performance as good or better than a larger number of tasks.
-In particular, the top subplot in column (b) suggests that two task models perform particularly well when given a relatively inexpensive training set of the CAT structure.
-Note that, in each model, only one secondary task is given access to T data.
+Note that we consider the same combination of sizes for the C, A, and T sets regardless of the total number of tasks, $N$.
+In the top row, $k=0$ implies that molecular configurations in the C and A sets are each assigned to exactly one secondary task, regardless of the total number of secondary tasks.
+Therefore, in this row, we can compare models implemented with different $N$ but nearly identical training budgets.
+The results show that for sufficiently large training sets, given a fixed budget, there is a benefit to dividing the budget across multiple secondary tasks rather than employing only one secondary task.
+For smaller datasets, choosing $N=2$ may lead to performance as good or better than a larger number of tasks.
+In particular, the top subplot in column (b) suggests that models where $N=2$ perform particularly well when given a relatively inexpensive training set of the CAT structure.
+Note that, in each model, only one of the secondary tasks is trained using T data: only $\mathcal{T}(\text{PBE})$ is nonempty.
 Thus, in this setting, the decrease in performance when new tasks are added likely comes from siphoning off part of the computational budget to train tasks which do not have access to T data.
-We can also consider whether it is useful to train on more than two predictions for the same molecular configuration.
-The middle row of Fig.
-6 shows cases where each task has access to training data for the same 50% of molecules in the C set as well as 50% of the molecules in the A set while the remaining molecules are distributed one per task, and the final row shows cases where all tasks have access to all data in the C and A sets.
-From these subplots, we conclude that there is a benefit to training our model on multiple secondary predictions for the same molecule.
+We can also consider whether it is useful to include the same molecular configuration in the training set of multiple secondary tasks.
+In the final row of Fig.
+6, $k=1$ implies that $\mathcal{C}(s_{n})\cup\mathcal{A}(s_{n})$ is the same set for all $n=1,\dots,N-1$.
+By contrast, in the middle row, to enforce $k=\frac{1}{2}$, we assign $50\%$ of the configurations in $C$ and $A$ to every secondary task while the remaining configurations in these sets are distributed between the secondary tasks so that each is assigned to exactly one secondary task.
+The improved performance of the $N=3$ and $N=5$ cases in the last row compared with the middle row suggest that there is a benefit to training our model on multiple secondary predictions for the same molecule.
 V Conclusion
@@
 Furthermore, use of the multitask framework can reduce the cost of generating entirely new training data by an order of magnitude and can also facilitate the opportunistic incorporation of existing datasets of heterogeneous quality into the training set.
-Our work is the first to explore a multitask GP regression model in a materials science setting.
-Significantly, the multitask framework is not restricted to the choice of GP inference, and future work can investigate training models with neural networks and other tools for regression.
+To our knowledge, our work is the first to explore an asymmetric multitask GP regression model in a materials science setting.
+Significantly, the multitask framework is not restricted to the choice of GP inference.
+Future work can preserve the same relationship between primary and secondary task submodels and investigate training each submodel with neural networks or other tools for regression.
 Unlike other machine learning methods for constructing surrogates from ab initio data, the multitask model does not require all data sources be ordered in a hierarchy of accuracy.
@@
 The data that support the findings of this study for the small organic molecules example are at https://doi.org/10.5281/zenodo.10215421.
-Appendix A Training Data Cost Model
+Appendix A Lengthscale and Variance Optimization
+Refer to caption Refer to caption
+(a) Primary hyperparameters (b) Secondary hyperparameters
+Figure 7: Water Trimer Case.
+Each point compares the test MAE of a GP model which uses the MLE variance to the test MAE of a model which uses a variance estimate obtained by minimizing MAE on the optimization set.
+Different data points are obtained by shuffling the optimization data, and color indicates the size of the optimization data set.
+In plot (a) we consider a GP model for the primary regression task while in plot (b) we consider a model $\delta_{s_{n}}$, as defined in (12).
+This section provides further detail on our procedure for hyperparameter optimization with a focus on computational tractability.
+We note that further improvements are possible and worth investigating in future work.
+Specifically, we discuss optimization for the test case explored in Subsections IV.2.1 and IV.2.2 to estimate the hyperparameters of the prior distributions for $f_{p}$ and $\delta_{s_{1}}$, given by (13) and (14).
+Both priors are constructed with the squared exponential kernel:
+$k(\bm{X},\bm{X}^{\prime})\ =\ v\ \exp\left(\frac{-(\bm{X}-\bm{X}^{\prime})^{T}(\bm{X}-\bm{X}^{\prime})}{2\ell}\right)$
+As discussed in IV.1.3, we optimize hyperparameters using a data set of $M=150$ molecular configurations.
+For all statistical models used in Subsections IV.2.1 and IV.2.2, the lengthscale, $\ell$, is obtained through a maximum likelihood procedure while the variance, $v$, is taken to be the value which minimizes MAE on the optimization set.
+The latter choice was made because, for the three body energy data, variance estimates produced by minimizing MAE achieve lower test error than the maximum likelihood estimate of variance.
+Fig.
+7 establishes that these choices of $M$ and variance estimate are not necessary for the multitask model to perform well; in fact, the model appears relatively robust to reasonable optimization procedures.
+Each color in Fig.
+7 represents a different optimization set size, $M$.
+We obtain multiple hyperparameter estimates for each $M$ by shuffling the optimization data set.
+For each shuffled realization of the data, we solve for estimates ^$\hat{\ell}_{MLE}$, $\hat{v}_{MLE}$, and $\hat{v}_{MAE}$.
+The plotted points compare MAE on a test data set of $320$ molecular configurations obtained using ^$\hat{\ell}_{MLE}$ and $\hat{v}_{MLE}$ to MAE for the same test data set obtained using ^$\hat{\ell}_{MLE}$ and $\hat{v}_{MAE}$.
+We can see that both approaches to estimating variance lead to good predictive performance, but use of $\hat{v}_{MAE}$ generally produces lower MAE than use of $\hat{v}_{MLE}$ for both $f_{p}$ and $\delta_{s_{1}}$.
+Fig.
+7 also demonstrates that the optimization data set size can be reduced without significant reductions in prediction accuracy though there is greater variance in prediction quality for $M\leq 45$.
+Note that given ^$\hat{\ell}_{MLE}$, the expression for $\hat{v}_{MLE}$ is closed form and $\hat{v}_{MAE}$ is found using ^$\hat{\ell}_{MLE}$.
+Thus, we only optimize for one hyperparameter at a time.
+If the optimization procedure reaches a threshold of 30 seconds, we halt the procedure, reshuffle the data, and restart.
+In the case that $M=150$, we require on the order of $10$ seconds on a laptop to successfully obtain MLE estimates.
+MAE estimates are found in $\sim 100$ seconds for the primary task and $\sim 40$ seconds for the secondary task.
+The reported amounts include any time spent on false starts which time out at $30s$.
+By comparison, a single DFT calculation for 3-b energy using a SCAN functional approximation requires $\sim 650$ seconds on a cluster.
+Thus, compared with data generation, hyperparameter optimization accounts for a small portion of the computational budget of our numerical experiments.
+Our results in Fig.
+7 suggest that we can construct successful statistical models even when the budget for generating the optimization data set is much smaller than in our work.
+For any $M$, the cost of generating the optimization set will be dominated by CCSD(T) computations which are necessary for all statistical models which we implement.
+Future work may fully evaluate the minimum cost required to produce hyperparameters which will achieve a specified threshold of accuracy and investigate strategies for constructing opportunistic optimization sets from existing data.
+Appendix B Training Data Cost Model
 In our numerical results, we estimate the training cost as
-model training $\text{model training cost}\ =\ \sum_{j}^{m}n_{j}\hat{c}_{j}$
-where $n_{j}$ is the number of training systems from the $j^{th}$ level of theory and $\hat{c}_{j}$ is an estimate of prediction cost for that level of theory.
-We produce one set of estimates for $\hat{c}_{1},\dots,\hat{c}_{m}$ for the water trimer example and a different set of estimates for the small organic molecules example.
+model training $\text{model training cost}\ =\ \sum_{\ell}^{L}n_{\ell}\hat{c}_{\ell}$
+where $n_{\ell}$ is the number of training systems from the $\ell^{th}$ level of theory and $\hat{c}_{\ell}$ is an estimate of prediction cost for that level of theory.
+We produce one set of estimates for $\hat{c}_{1},\dots,\hat{c}_{L}$ for the water trimer example and a different set of estimates for the small organic molecules example.
 For the water trimers example, we randomly selected ten molecular configurations from our data set as the basis for our cost model.
@@
 All computations for the water cost model where performed on physically the same Intel Xeon Platinum 8260 node of MIT SuperCloud (https://supercloud.mit.edu/systems-and-software).
-All computations for a given molecular configuration where performed on the same basis set subasis set sucore.
+All computations for a given molecular configuration where performed on the same core.
 For this case study, we estimate that a CCSD(T) calculation incurs $37$ times the cost of a DFT calculation based on the average prediction time of these methods on our example molecular configurations.
-Several configurations in the small organic molecules data set are larger than the water trimer configurations.
-The median number of electrons in this data set is $40$, the minimum is $14$, and the maximum is $50$.
-Due to the expense of ionization potential calculations, we base our cost model on energy calculations for ten randomly sampled configurations with $36$ electrons.
-Note that CCSD(T) scales as $N^{7}$ where $N$ is the number of electrons in a system while DFT scales as $N^{3}$.
+Several systems in the small organic molecules data set contain more atoms than the water trimer configurations.
+The median number of electrons per system in the organic molecules set is $40$, the minimum is $14$, and the maximum is $50$.
+We estimate the cost of level of theory $\ell$ as the computational time required to obtain the total energy at this level of theory averaged over ten randomly sampled configurations with $36$ electrons.
+Note that this estimate is conservative: CCSD(T) scales as $N_{e}^{7}$ where $N_{e}$ is the number of electrons in a system while DFT scales as $N_{e}^{3}$.
 Therefore, an estimate of the ratio between the cost of CCSD(T) and the cost of DFT based on $36$ electrons will be smaller than the actual ratio for most molecular configurations in our data set.
-Furthermore, the prediction of ionization potential was performed using ${\Delta}CC$ (delta coupled cluster).
+Furthermore, our prediction of the ionization potential was performed using ${\Delta}CC$ (delta coupled cluster), which requires multiple CCSD(T) calculations to reach an IP prediction.
 Thus, while our reported results (Fig.
 6) estimate that CCSD(T) calculations demand $244$ times the expense of DFT, the true ratio between the cost of CCSD(T) and DFT is even larger.
-Appendix B SOAP parameters
+Consequently the reported gains in efficiency for the multitask method are expected to be a lower bound on the true improvement.
+Appendix C SOAP parameters
 To fix reasonable values for parameters required by SOAP ($r_{cut}$, $\sigma_{atom}$, $n_{max}$, $l_{max}$), we refer to established conventions as well as experimentation.
@@
 For water trimers, we use $r_{cut}=10\ $\mathrm{\SIUnitSymbolAngstrom}$$ to capture all three component water molecules.
-Appendix C Global Features
+Appendix D Global Features
 Figure 9: Correlation between different methods of constructing global features when calculating the distance between molecule pairs. De et al. (2016)
@@
 These two strategies are also both positively correlated with the “Average by Species” approach which produces an averaged feature for each element in the system and inserts isolated atoms to represent elements not included in the system, but these correlations are not as strong as that between the REMatch and totally averaged features.
-We may see this result because the REMatch and totally averaged features share the same dimension, $n$, whereas features that are constructed by species specific averaging have dimension of $2n$ to $4n$ in the cases we considered.
+We may see this result because the REMatch and totally averaged features share the same dimension, $n$, whereas features that are constructed by species-specific averaging have dimension of $2n$ to $4n$ in the cases we considered.
 Because averaging features is much more computationally efficient than constructing REMatch features, we use this approach in our numerical experiments.